### v1

# Enhanced Rényi Entropy-Based Post-Quantum Key Agreement with Provable Security and Information-Theoretic Guarantees

Abstract
This paper presents an enhanced post-quantum key agreement protocol based on Rényi entropy, addressing vulnerabilities in the original construction while preserving information-theoretic security properties. We develop a theoretical framework leveraging entropy-preserving operations and secret-shared verification to achieve provable security against quantum adversaries. Through entropy amplification techniques and quantum-resistant commitments, the protocol establishes $2^{128}$ quantum security guarantees under the quantum random oracle model. Key innovations include a confidentiality-preserving verification mechanism using distributed polynomial commitments, tightened min-entropy bounds with guaranteed non-negativity, and composable security proofs in the quantum universal composability framework. Unlike computational approaches, our method provides information-theoretic security without hardness assumptions while maintaining polynomial complexity. Theoretical analysis demonstrates resilience against known quantum attack vectors, including Grover-accelerated brute force and quantum memory attacks. The protocol achieves parameterization for 128-bit quantum security with efficient $\mathcal{O}(n^{2})$ communication complexity. Extensions to secure multiparty computation and quantum network applications are established, providing a foundation for long-term cryptographic security. All security claims are derived from mathematical proofs; this theoretical work presents no experimental validation.
Index Terms:
Rényi entropy, post-quantum cryptography, key agreement, information-theoretic security, quantum security, entropy preservation, secret sharing, quantum random oracle model.
1 Introduction
The anticipated development of large-scale quantum computing represents a significant challenge to contemporary cryptographic infrastructure [31, 20, 14, 1]. Shor’s polynomial-time quantum algorithm for integer factorization and discrete logarithms [1] compromises the security of widely-deployed asymmetric cryptosystems including RSA, ECC, and Diffie-Hellman key exchange. Grover’s quadratic speedup for unstructured search [2] reduces the effective security strength of symmetric primitives, necessitating key size increases [27]. This quantum threat landscape has catalyzed post-quantum cryptography (PQC) development, with NIST standardizing lattice-based, code-based, and multivariate schemes [20, 9]. However, these approaches rely on computational hardness assumptions that may be vulnerable to algorithmic advances [26, 28, 10].
1-A Research Context and Motivation
Information-theoretic cryptography offers an alternative paradigm with unconditional security guarantees that persist against quantum adversaries possessing unbounded computational resources [21, 15]. While quantum key distribution (QKD) provides information-theoretic security [25, 12], it requires specialized quantum communication hardware and authenticated classical channels. Classical information-theoretic solutions based on secret sharing [4] and physical unclonable functions exist but typically require pre-shared keys or lack quantum resistance [17].
The original Rényi entropy-based key agreement protocol [32] demonstrated theoretical promise but contained vulnerabilities: (1) input exposure during broadcast enabling key compromise, (2) potentially negative entropy bounds violating security requirements, and (3) insufficient protection against quantum-specific attacks. This work addresses these limitations through three key innovations:
  1. 1.
    Quantum-resistant distributed verification (Section 3.2): Novel polynomial commitment scheme based on Shamir’s secret sharing [4] with information-theoretic confidentiality
  2. 2.
    Entropy amplification with guaranteed positivity (Section 4.1): Rigorous min-entropy bounds for XOR composition with non-negativity constraints
  3. 3.
    Composable security framework (Section 5.3): Formal security proofs in quantum universal composability model [5, 6]
Research Purpose: To establish a theoretical framework for information-theoretically secure key agreement resistant to quantum attacks, providing a foundation for long-term cryptographic security without reliance on computational hardness assumptions.
The theoretical foundation of our approach rests on three interconnected pillars of quantum information theory: (1) Quantum entropy preservation - ensuring min-entropy bounds hold against quantum adversaries with side information; (2) Distributed verification - enabling secure consistency checks without exposing sensitive inputs; (3) Composable security - providing security guarantees under quantum composition.
Theorem 1 (Quantum Entropy Preservation Bound).
For $n$ independent entropy sources $s_{i}$ with $H_{\infty}(s_{i})\geq\gamma$ and quantum adversary holding quantum side information $\rho_{E}$, the combined secret $S=\bigoplus_{i=1}^{n}s_{i}$ satisfies:
$S_{\infty}(S|\rho_{E})\geq n\gamma-(n-1)m-S_{0}(\rho_{E})$ (1)
where $S_{0}(\rho_{E})$ is the quantum max-entropy of $\rho_{E}$.
Theorem 1 establishes the core security foundation, demonstrating that the XOR combination preserves min-entropy even when adversaries possess quantum side information. The $S_{0}(E)$ term quantifies the security degradation from quantum memory attacks, which we mitigate through parameter optimization.
1-B Key Contributions
This paper makes seven significant advances in post-quantum cryptography:
1. Confidentiality-Preserving Verification: A secret-shared verification mechanism using distributed polynomial commitments that prevents input exposure while enabling secure entropy verification, addressing a vulnerability in prior constructions [32].
2. Optimal Entropy Bounds: Mathematically rigorous min-entropy preservation theorems with guaranteed non-negative bounds and parameterization strategies for quantum security.
3. Quantum Universal Composability: Composable security proof for information-theoretic key agreement in the quantum universal composability (QUC) model [5, 6], demonstrating secure realization of ideal key functionality.
4. Quantum Attack Resistance: Formal analysis against known quantum attack vectors including Grover-accelerated brute force, quantum collision search [7], and quantum memory attacks [21], with quantifiable security bounds.
5. Entropy Amplification Framework: Generalized framework for multi-party entropy amplification using R’enyi entropy measures [3], providing exponential security scaling against quantum adversaries.
6. Hybrid Security Extension: Integration with quantum key distribution (QKD) [25] that enhances security against active adversaries while preserving information-theoretic guarantees.
7. Secure Computation Extension: Secure extension to general secure multiparty computation [24, 17] for linear functions, enabling privacy-preserving applications.
These contributions establish a paradigm for quantum-resistant cryptography based on information-theoretic principles, with applications in secure multiparty computation and quantum networks.
1-C Quantum Entropic Framework
The foundation of our approach rests upon the rigorous quantification of uncertainty in quantum systems. Unlike classical entropy measures, quantum Rényi entropy (9) captures the fundamental limits of information extraction under quantum mechanical constraints. This becomes crucial when adversaries possess quantum memory capable of storing superposition states for delayed measurement [21]. Our framework explicitly addresses this quantum advantage by establishing composable entropy bounds through:
  • Quantum-proof min-entropy extraction from non-uniform sources
  • Entropic uncertainty relations under quantum side information
  • Tight bounds on quantum guessing probability via $S_{\infty}(\rho)$
This theoretical foundation enables security guarantees that persist even when adversaries exploit quantum coherence and entanglement.
2 Background and Theoretical Foundations
2-A Quantum Computing Threat Model
Quantum computation utilizes state superposition and quantum correlation effects to achieve computational advantages for specific problems [14]. We formalize the quantum adversary model:
Definition 1 (Quantum Polynomial-Time Adversary).
A quantum adversary $\mathscr{A}$ is a polynomial-time quantum algorithm with:
  • Quantum random oracle access to hash functions
  • Quantum memory bounded by $Q=2^{\mathcal{O}(\kappa)}$ qubits
  • Capability to corrupt up to t1t-1 parties
  • Ability to perform superposition queries to oracles
  • Adaptive measurement strategies [21, 30]
Threat Analysis:
  • Shor’s Algorithm: Factors integers in $\mathcal{O}((\log N)^{3})$ time [1], compromising RSA, ECC, and Diffie-Hellman
  • Grover’s Algorithm: Solves unstructured search in $\mathcal{O}(\sqrt{N})$ time [2], reducing symmetric key security
  • Brassard-Høyer-Tapp (BHT): Quantum collision finding in $\mathcal{O}(2^{m/3})$ time [7]
  • Quantum Memory Attacks: Exploit coherent state persistence in quantum storage [21, 23]
  • Quantum Rewinding: Subvert classical proof systems [31]
Research Significance: Our protocol design counters these threats through entropy amplification and information-theoretic security [30, 16], ensuring security against quantum adversaries.
2-B Rényi Entropy Framework
Rényi entropy [3] provides a parametric family of entropy measures essential for cryptographic security analysis. Its operational significance in quantum cryptography stems from connections to guessing probabilities [18]:
Definition 2 (Rényi Entropy).
For discrete random variable $X\sim P_{X}$ over $\mathcal{X}$, Rényi entropy of order $\alpha$ is:
$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\left(\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}\right)$ (2)
for $\alpha>0$, $\alpha\neq 1$.
Cryptographically significant special cases:
$\lim_{\alpha\to 1}H_{\alpha}(X)=H(X)=-\sum_{x}P_{X}(x)\log P_{X}(x)$ (Shannon) (3)
$H_{2}(X)=-\log\left(\sum_{x}P_{X}(x)^{2}\right)$ (Collision) (4)
$\lim_{\alpha\to\infty}H_{\alpha}(X)=H_{\infty}(X)=-\log\max_{x}P_{X}(x)$ (Min-entropy) (5)
Operational Significance [18, 15, 19]:
  • Min-entropy: $H_{\infty}(X)=-\log P_{\text{guess}}(X)$ where $P_{\text{guess}}$ is optimal guessing probability
  • Collision entropy: $H_{2}(X)=-\log P_{\text{coll}}(X)$ where $P_{\text{coll}}$ is collision probability
  • Quantum advantage: Min-entropy bounds quantum guessing probability under side information
Fundamental properties enabling our security proofs:
Lemma 1 (Entropy Transformation).
For independent $X,Y$ and deterministic function $f$:
$H_{\infty}(f(X))$ $\geq H_{\infty}(X)-\log|\text{range}(f)|$ (6)
$H_{\infty}(X\oplus Y)$ $\geq\max(H_{\infty}(X),H_{\infty}(Y))$ (7)
$H_{\infty}(X\oplus Y)$ $\geq H_{\infty}(X)+H_{\infty}(Y)-\log|\mathcal{X}|$ (8)
Proof.
Follows from probability bound $\max_{z}\Pr[f(X)=z]\geq\frac{\max_{x}\Pr[X=x]}{|\text{range}(f)|}$ and convolution properties. Detailed proof in [13]. ∎
2-C Quantum Information Theoretic Foundations
Definition 3 (Classical vs. Quantum Entropy Notation).
Throughout this paper:
  • $H_{\alpha}(X)$ denotes classical Rényi entropy for random variable $X$
  • $S_{\alpha}(\rho)$ denotes quantum Rényi entropy for density operator $\rho$
  • $H_{\infty}(X)$ and $S_{\infty}(\rho)$ specifically denote min-entropy
For quantum systems, security analysis requires quantum entropy measures [19]:
Definition 4 (Quantum Rényi Entropy).
For density operator $\rho$, quantum Rényi entropy is:
$S_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\operatorname{tr}(\rho^{\alpha})$ (9)
Quantum Security Bounds [15, 16, 29, 23]:
  • Quantum min-entropy: $S_{\infty}(\rho)=-\log\lambda_{\max}(\rho)$ bounds state guessing probability
  • Quantum collision entropy: $S_{2}(\rho)=-\log\operatorname{tr}(\rho^{2})$ quantifies state distinguishability
  • Data processing inequality: $S_{\alpha}(\mathscr{E}(\rho))\geq S_{\alpha}(\rho)$ for quantum channels $\mathscr{E}$
  • Entropy accumulation: $S_{\infty}(\rho_{X^{n}E})\geq\sum_{i=1}^{n}S_{\infty}(\rho_{X_{i}E})-c\sqrt{n}$ [23]
Research Significance: Our security proofs leverage these quantum entropy measures to establish composable security against quantum adversaries.
2-D Distributed Verification via Secret Sharing
Shamir’s secret sharing [4] provides the foundation for our confidentiality-preserving verification:
Theorem 2 (Shamir’s Secret Sharing).
For secret $s\in\mathbb{F}_{q}$, threshold $t$, and $n$ parties, choose random polynomial:
$f(x)=s+a_{1}x+\cdots+a_{t-1}x^{t-1}\in\mathbb{F}_{q}[x]$ (10)
Distribute shares $s_{i}=f(i)$. Then:
  • Any $t$ shares reconstruct $s$ via Lagrange interpolation
  • Any t1t-1 shares reveal zero information about $s$
We enhance this scheme with verifiable features [24]:
  • Polynomial commitments: Binding cryptographic commitments to coefficients
  • Distributed verification: Consistency checks without reconstruction
  • Information-theoretic confidentiality: Guaranteed by secret sharing properties
2-E Symbol Table
TABLE 1: Mathematical Symbols and Definitions
Symbol Meaning
$n$ Number of protocol participants
$\kappa$ Security parameter (bit strength)
$m$ Bit-length of secret shares
$\gamma$ Min-entropy threshold per share
$t$ Reconstruction threshold ($t=\lfloor n/2\rfloor+1$)
$\delta$ Entropy estimation accuracy
$H_{\alpha}(X)$ Classical Rényi entropy of order $\alpha$
$S_{\alpha}(\rho)$ Quantum Rényi entropy
$H_{\infty}(X)$ Min-entropy: $\log\max_{x}P_{X}(x)$
$\hat{H}_{i}$ Estimated min-entropy of $s_{i}$
$s_{i}$ Secret share generated by party $i$
$c_{i}$ Cryptographic commitment: $\mathcal{H}(s_{i}\|\hat{H}_{i})$
$f_{i}(x)$ Sharing polynomial for $s_{i}$
$sh_{ij}$ Share of $s_{i}$ distributed to $P_{j}$
$\mathbb{F}_{2^{m}}$ Finite field with $2^{m}$ elements
$\mathcal{H}$ Quantum random oracle
$\oplus$ Bitwise XOR operation
$Q$ Adversary quantum memory bound ($Q=2^{\kappa/2}$ qubits)
$\mathscr{A}$ Quantum polynomial-time adversary
$\mathcal{F}_{KA}$ Ideal key agreement functionality
$K$ Established secret key
$\lambda$ Commitment security parameter ($\lambda=2\kappa$)
$\epsilon$ Security error bound ($\epsilon=2^{-40}$)
3 Enhanced Protocol Design
3-A System Model and Threat Analysis
Consider a network of $n$ parties $\mathcal{P}=\{P_{1},\dots,P_{n}\}$ establishing shared secret key $K\in\{0,1\}^{\kappa}$.
Adversarial Capabilities [6, 31, 21]:
  • Quantum polynomial-time computation
  • Quantum random oracle access to $\mathcal{H}$
  • Quantum memory bounded by $Q=2^{\kappa/2}$ qubits
  • Adaptive corruption of up to t1t-1 parties
  • Control over communication channels
  • Superposition queries and delayed measurement
  • Quantum rewinding attacks
Security Assumptions:
  • Authenticated secure channels between parties
  • Reliable entropy sources with $\delta$-accurate min-entropy estimation
  • Quantum-resistant hash function $\mathcal{H}$ modeled as QROM
  • Honest majority ($t=\lfloor n/2\rfloor+1$)
  • Entropy independence across parties
Design Rationale: This model balances practical quantum threats with theoretical tractability, enabling security proofs while capturing essential quantum capabilities.
3-B Confidentiality-Preserving Verification Mechanism
The original vulnerability stemmed from broadcasting $s_{i}$, allowing adversaries to compute $K=\mathcal{H}(\oplus_{i}s_{i})$. Our solution integrates secret sharing with cryptographic commitments:
Definition 5 (Distributed Polynomial Commitment).
For secret $s_{i}$, generate random polynomial:
$f_{i}(x)=s_{i}+\sum_{k=1}^{t-1}a_{k}x^{k}\in\mathbb{F}_{2^{m}}[x]$ (11)
with $a_{k}\xleftarrow{\$}\{0,1\}^{m}$. The commitment is $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ where $\hat{H}_{i}\approx H_{\infty}(s_{i})$.
Verification Protocol:
  1. 1.
    Each party $P_{j}$ receives share $sh_{ij}=f_{i}(j)$
  2. 2.
    $P_{j}$ collects $t$ shares $\{sh_{ik}\}_{k\in S}$
  3. 3.
    Reconstructs $\tilde{s}_{i}$ via Lagrange interpolation
  4. 4.
    Verifies $c_{i}=\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})$
  5. 5.
    Verifies $H_{\infty}(\tilde{s}_{i})\geq\gamma-\delta$
Security Properties [24, 4, 6]:
  • Confidentiality: <t<t colluding parties learn nothing about $s_{i}$
  • Binding: Computational binding under QROM
  • Verifiability: Algebraic structure enables consistency checks
  • Quantum resistance: Security preserved under superposition attacks
The distributed verification mechanism solves the input exposure vulnerability through a novel application of polynomial commitments over secret shares. Unlike standard commitment schemes, our approach preserves information-theoretic confidentiality while providing computational binding under quantum attacks. This dual-security property is achieved through the algebraic structure of Shamir’s secret sharing, where verification occurs locally on shares without reconstruction.
Theorem 3 (Verification Security).
For any QPT adversary $\mathscr{A}$ attempting to submit invalid share $s^{\prime}\neq s$:
$\Pr[\text{successful verification}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}$ (12)
where $q$ is the number of quantum queries to $\mathcal{H}$.
Theorem 3 quantifies the security of our verification mechanism against quantum adversaries. The $\frac{q^{2}}{2^{m}}$ term bounds quantum collision attacks on the commitment scheme, while $\frac{1}{2^{\gamma-\delta}}$ represents the probability of guessing a valid high-entropy secret. By setting $m=3\kappa$ and $\gamma>\kappa$, both terms become negligible for $\kappa\geq 128$.
3-C Key Agreement Protocol
The complete protocol operates in four phases with comprehensive quantum resistance:
Algorithm 1 Enhanced Rényi Key Agreement Protocol
0: Security parameter $\kappa$, min-entropy threshold $\gamma$, parties $n$, estimation accuracy $\delta$
0: Shared secret key $K$
 Set $t=\lceil n/2\rceil+1$ {Honest majority threshold}
Phase 1: Initialization
for each party $P_{i}\in\mathcal{P}$ do
  Sample $s_{i}\leftarrow\{0,1\}^{m}$ with $H_{\infty}(s_{i})\geq\gamma$ {High-entropy secret}
  Estimate $\hat{H}_{i}$ such that $|H_{\infty}(s_{i})-\hat{H}_{i}|\leq\delta$ {Min-entropy estimation}
  Compute commitment $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ {QROM-based binding}
  Broadcast $c_{i}$ to all parties {Public commitment}
end for
Phase 2: Share Distribution
for each $P_{i}\in\mathcal{P}$ do
  Generate random polynomial $f_{i}(x)=s_{i}+\sum_{k=1}^{t-1}a_{k}x^{k}$, $a_{k}\xleftarrow{\$}\{0,1\}^{m}$
  for each $P_{j}\in\mathcal{P}\setminus\{P_{i}\}$ do
   Compute share $sh_{ij}=f_{i}(j)$
   Securely send $(i,sh_{ij})$ to $P_{j}$ via authenticated channel
  end for
end for
Phase 3: Verification
for each $P_{j}\in\mathcal{P}$ do
  for each $P_{i}\in\mathcal{P}\setminus\{P_{j}\}$ do
   Collect $\geq t$ valid shares $\{sh_{ik}\}_{k\in S}$ for $|S|\geq t$
   Reconstruct $\tilde{s}_{i}=\sum_{k\in S}sh_{ik}\cdot L_{k}(0)$ {Lagrange interpolation}
   Verify $\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})=c_{i}$ {Commitment consistency}
   Verify $H_{\infty}(\tilde{s}_{i})\geq\gamma-\delta$ {Entropy threshold}
   if any verification fails then
    Abort protocol and output $\bot$
   end if
  end for
end for
Phase 4: Key Derivation
 Each $P_{i}$ reveals $s_{i}$ to all parties via authenticated channels {Safe after verification}
 Compute $S=\bigoplus_{i=1}^{n}s_{i}$ {Entropy-preserving combination}
 Compute $K=\mathcal{H}(S)$ {Quantum-secure randomness extraction}
return  $K$
Security Justification for Revealing Phase: After successful verification, revealing $s_{i}$ through authenticated channels is secure because: (1) Verification ensures each $s_{i}$ has min-entropy $\geq\gamma-\delta$, (2) The final key $K$ is derived from XOR of all $s_{i}$ followed by hashing, (3) Adversaries cannot alter $s_{i}$ due to authenticated channels, and (4) Commitment binding prevents submission of different values.
Protocol Properties:
  • Communication Complexity: $\mathcal{O}(n^{2}m)$ bits
  • Computational Complexity: $\mathcal{O}(n^{2})$ field operations
  • Round Complexity: 3 rounds (broadcast, share exchange, key derivation)
  • Fault Tolerance: Resilient to <t<t malicious parties
3-D Entropy Verification Theoretical Basis
Theoretical Challenge: Exact min-entropy computation requires complete knowledge of the underlying probability distribution, which is generally infeasible.
Resolution Framework [18, 13, 15]:
  1. 1.
    Assume parties have access to entropy sources with certified min-entropy bounds
  2. 2.
    Utilize statistical estimation techniques:
    $\hat{H}_{\infty}=-\log\left(\max_{x\in\mathcal{S}}\hat{P}(x)\right)\pm\delta$ (13)
    for empirical distribution $\hat{P}$ over sample $\mathcal{S}$ of size $N\geq 2^{2\gamma}/\epsilon^{2}$ to achieve $\delta$-accuracy with failure probability $\epsilon$
  3. 3.
    Incorporate estimation error $\delta$ into security margins via $\gamma-\delta$
  4. 4.
    Quantum extension: Use quantum-proof estimators [15]
Security Implications: The verification condition $H_{\infty}(\tilde{s}_{i})\geq\gamma-\delta$ ensures security even with bounded estimation error. The $\delta$ parameter must be conservatively chosen to account for statistical uncertainty.
3-E Quantum Security Design Principles
The protocol architecture incorporates three fundamental quantum-resistant mechanisms:
  1. 1.
    Superposition Resistance: Polynomial commitments use injective encoding $\mathcal{H}(s_{i}\parallel\hat{H}_{i})$ to prevent quantum ambiguity attacks. The concatenation ensures $\forall s_{i}\neq s_{j},\mathcal{H}(s_{i}\parallel\cdot)\cap\mathcal{H}(s_{j}\parallel\cdot)=\emptyset$ with probability $1-\operatorname{negl}(\lambda)$ under QROM.
  2. 2.
[…]
    $\Delta(\rho_{SE},\rho_{S}\otimes\rho_{E})\leq 2^{-H_{\min}(S|E)}$ (14)
    where $\Delta$ is trace distance.
  3. 3.
    Quantum Rewinding Protection: The verification phase forces sequential measurement through:
    Commit $\rightarrow\text{Share}\rightarrow\text{Verify}$
    $\Downarrow$
    No-cloning theorem enforced
    Adversaries cannot maintain superposition beyond verification due to the measurement requirement in Step 3 of Algorithm 1.
4 Enhanced Theoretical Analysis
4-A Entropy Amplification Theorem
The security of key derivation relies on min-entropy preservation during XOR combination:
Theorem 4 (Min-Entropy Preservation).
For independent random variables $X_{1},\dots,X_{n}$ over $\{0,1\}^{m}$ with $H_{\infty}(X_{i})\geq\gamma$, the XOR sum $S=\bigoplus_{i=1}^{n}X_{i}$ satisfies:
$H_{\infty}(S)\geq\max\left(0,n\gamma-(n-1)m\right)\geq\kappa$ (15)
Equality holds when $X_{i}$ are uniform and independent.
Proof.
By induction on $n$.
Base Case ($n=2$): For independent $X,Y$ over $\{0,1\}^{m}$:
$\max_{z}\Pr[X\oplus Y=z]$ $=\max_{z}\sum_{x}\Pr[X=x]\Pr[Y=z\oplus x]$
$\leq\max_{z}\sum_{x}\Pr[X=x]\cdot\max_{y^{\prime}}\Pr[Y=y^{\prime}]$
$=\max_{y^{\prime}}\Pr[Y=y^{\prime}]\cdot\sum_{x}\Pr[X=x]$
$=2^{-H_{\infty}(Y)}$ (16)
However, a tighter bound derives from convolution properties. For finite field $\mathbb{F}_{2^{m}}$, we have:
$\max_{z}\Pr[X\oplus Y=z]\leq\min\left(2^{-H_{\infty}(X)},2^{-H_{\infty}(Y)}\right)\cdot|\mathcal{X}|\cdot\alpha$ (17)
where $\alpha=\sup_{x,y}\Pr[Y=y|X=x]$. Under independence $\alpha=2^{-H_{\infty}(Y)}$, yielding:
$H_{\infty}(X\oplus Y)\geq\max\left(H_{\infty}(X),H_{\infty}(Y)\right)+\log|\mathcal{X}|-\beta$ (18)
with $\beta=\log(1+\operatorname{corr}(X,Y))$. For independent variables $\beta=0$, recovering Equation (19). Full derivation follows Vadhan [13] (Lemma 6.21). Similarly, $\max_{z}\Pr[X\oplus Y=z]\leq 2^{-H_{\infty}(X)}$. Thus:
$H_{\infty}(X\oplus Y)\geq\max(H_{\infty}(X),H_{\infty}(Y))$
The tighter bound follows from the convolution inequality:
$H_{\infty}(X\oplus Y)\geq H_{\infty}(X)+H_{\infty}(Y)-\log|\mathcal{X}|.$ (19)
This inequality holds because the XOR operation reduces the maximum probability by at most $\log|\mathcal{X}|$ due to the discrete nature of the alphabet. For any $z$, we have:
$\Pr[X\oplus Y=z]$
$\leq\max_{x}\Pr[X=x]\cdot\max_{y}\Pr[Y=y]$
$\leq 2^{-H_{\infty}(X)}\cdot 2^{-H_{\infty}(Y)}$ (20)
Taking the logarithm base 2, we get:
$H_{\infty}(X\oplus Y)$ $\geq-\log\max_{z}\Pr[X\oplus Y=z]$
$\geq-\log(2^{-H_{\infty}(X)}\cdot 2^{-H_{\infty}(Y)})$
$=H_{\infty}(X)+H_{\infty}(Y).$ (21)
However, due to the finite alphabet size, the bound is slightly weaker:
$H_{\infty}(X\oplus Y)\geq H_{\infty}(X)+H_{\infty}(Y)-\log|\mathcal{X}|$ (22)
This completes the proof of the convolution inequality.
Inductive Step: Assume true for n1n-1 variables. Let $S_{n-1}=\bigoplus_{i=1}^{n-1}X_{i}$. Then:
$H_{\infty}(S)$ $=H_{\infty}(S_{n-1}\oplus X_{n})$
$\geq H_{\infty}(S_{n-1})+H_{\infty}(X_{n})-\log|\mathcal{X}|$
$\geq\left[(n-1)\gamma-(n-2)\log|\mathcal{X}|\right]+\gamma-\log|\mathcal{X}|$
$=n\gamma-(n-1)\log|\mathcal{X}|.$ (23)
The $\max(0,\cdot)$ ensures non-negativity when $n\gamma<(n-1)\log|\mathcal{X}|$. ∎
Quantum Extension [15, 21, 29]: For quantum side information $E$, by the chain rule:
$S_{\infty}(S|E)\geq n\gamma-(n-1)m-S_{0}(E)$ (24)
where $S_{0}(E)$ is the quantum max-entropy of $E$. Under bounded quantum memory $Q=2^{q}$ qubits, $S_{0}(E)\leq q$, yielding:
$S_{\infty}(S|E)\geq n\gamma-(n-1)m-q$
Lemma 2 (Quantum Chain Rule for Min-Entropy).
For quantum state $\rho_{X_{1}\cdots X_{n}E}$ and $S=\bigoplus_{i=1}^{n}X_{i}$:
$S_{\infty}(S|E)\geq\sum_{i=1}^{n}S_{\infty}(X_{i}|X_{1}\cdots X_{i-1}E)-(n-1)\log|\mathcal{X}|.$ (25)
Proof.
By induction and quantum data processing inequality. For $n=2$:
$S_{\infty}(X_{1}\oplus X_{2}|E)$ $\geq S_{\infty}(X_{1}\oplus X_{2}|X_{1},E)$
$=S_{\infty}(X_{2}|X_{1},E)$
$\geq S_{\infty}(X_{2}|E)+S_{\infty}(X_{1}|E)-\log|\mathcal{X}|.$
The inductive step follows from recursive application. Critical observation: conditioning on $X_{1}$ does not decrease entropy when $X_{1}$ is independent of $X_{2}E$. ∎
Corollary 1.
Under pairwise independence: $S_{\infty}(S|E)\geq\sum S_{\infty}(X_{i}|E)-(n-1)\log|\mathcal{X}|$.
4-B Security Parameterization Framework
Theorem 5 (Security Parameterization).
The protocol achieves $\kappa$-bit quantum security if:
$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$ (26)
for security parameter $\epsilon$, where $\delta$ accounts for entropy estimation error.
Proof.
From Theorem 15 and entropy estimation, $H_{\infty}(S)\geq n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$. By the quantum leftover hash lemma [15], for quantum-secure extractor $\mathcal{H}$:
$\delta(\rho_{KE},\tau_{K}\otimes\rho_{E})\leq\epsilon$ (27)
where $\delta$ is trace distance, $\tau_{K}$ uniform key, and $\rho_{E}$ quantum side information. Thus $K$ is $\epsilon$-close to uniform independent of $E$. ∎
Parameter Optimization: Solve:
$\min_{m,\gamma}$ $\quad n\gamma+(n-1)m$
s.t. $\quad n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$
$\quad m\geq\max(3\kappa,\gamma).$ (28)
$\quad\gamma\geq\gamma_{\min}$
Optimal solution: Set $m=3\kappa$ for BHT resistance, then $\gamma=\frac{\kappa+\log(1/\epsilon)+n\delta+(n-1)m}{n}$.
4-C Composable Security Framework
We prove security in the quantum universal composability (QUC) model [5, 6, 31]:
Theorem 6 (Composable Security).
The protocol $\Pi$ securely realizes the ideal key agreement functionality $\mathcal{F}_{KA}$ in the $(\mathcal{H},\mathcal{AUTH})$-hybrid model against QPT adversaries.
Proof.
Ideal Functionality $\mathcal{F}_{KA}$:
  • Upon receiving $(\text{init})$ from all honest parties, output $K\xleftarrow{\$}\{0,1\}^{\kappa}$ to all parties
  • Adversary learns nothing beyond protocol messages
Simulator Construction: For QPT adversary $\mathcal{A}$, construct simulator $\mathcal{S}$:
  1. 1.
    Commit phase: Generate random $c_{i}^{\prime}\xleftarrow{\$}\{0,1\}^{\lambda}$ without knowing $s_{i}$
  2. 2.
    Share phase: Simulate shares using random polynomials consistent with $c_{i}^{\prime}$
  3. 3.
    Verification phase: Program random oracle $\mathcal{H}$ to satisfy $\mathcal{H}(s_{i}^{\prime}\|\hat{H}_{i}^{\prime})=c_{i}^{\prime}$ for extracted $s_{i}^{\prime}$
  4. 4.
    Extraction: Monitor $\mathcal{A}$’s oracle queries to extract inputs via quantum rewinding [6]
  5. 5.
    Key derivation: Output $K$ consistent with extracted inputs
Indistinguishability: For any environment $\mathcal{Z}$, the distinguishing advantage is bounded by:
$\left|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{KA},\mathcal{S},\mathcal{Z}}=1]\right|\leq\mathsf{negl}(\kappa)$ (29)
This follows from:
  • Indistinguishability of commitments under QROM
  • Information-theoretic secrecy of shares (<t<t parties)
  • Programmability of random oracle [6]
  • Entropy preservation ensuring key uniformity
  • Quantum rewinding security [31]
5 Security Analysis
5-A Passive Security Against Quantum Eavesdroppers
Theorem 7 (Passive Security).
Under QROM, for any QPT passive adversary $\mathscr{A}$, the distinguishing advantage satisfies:
$\left|\Pr[\mathscr{A}(K)=1]-\Pr[\mathscr{A}(U)=1]\right|\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\mathsf{negl}(\lambda)$ (30)
where $U$ is uniform random, $q$ is number of quantum queries.
Proof.
Adversary advantage decomposes as:
$\text{Adv}(\mathscr{A})$ $\leq\delta(\text{real protocol},\text{ideal protocol})$
$+\delta(\text{ideal protocol},\text{uniform}).$ (31)
The first term is negligible by composability (Theorem 3). The second term is bounded by $2^{-\kappa}$ via quantum leftover hash lemma (27). The $\frac{q^{2}}{2^{m}}$ term bounds quantum collision probability in commitment scheme. ∎
5-B Active Adversary Resistance
Theorem 8 (Active Security).
With authenticated channels, for any QPT active adversary $\mathscr{A}$:
$\Pr[\mathscr{A}\text{ wins}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}+\frac{n^{2}}{2^{\lambda}}+\mathsf{negl}(\kappa)$ (32)
where win conditions include: (1) forcing acceptance of invalid shares, (2) distinguishing $K$ from random, or (3) causing honest parties to output different keys.
Proof.
Adversary wins by succeeding in at least one of:
  1. 1.
    Auth forgery: Forge authentication on $\geq t$ shares, probability $\mathsf{negl}(\lambda)$
  2. 2.
    Commitment collision: Find $s_{j}^{\prime}\neq s_{j}$ with $\mathcal{H}(s_{j}^{\prime}\|\hat{H}_{j}^{\prime})=c_{j}$, probability $\mathcal{O}(q^{2}/2^{m})$ by QROM collision resistance
  3. 3.
    Entropy fraud: Satisfy $H_{\infty}(s_{j}^{\prime})\geq\gamma-\delta$ for $s_{j}^{\prime}\neq s_{j}$, probability $\leq 2^{-(\gamma-\delta)}$
  4. 4.
    Share manipulation: Alter $\geq t$ shares without detection, prevented by binding property
Union bound gives the result. The $\frac{n^{2}}{2^{\lambda}}$ term accounts for commitment forgeries across all parties. ∎
5-C Quantum Attack Resilience Analysis
Grover’s Algorithm Resistance: The search space for $S=\oplus s_{i}$ has size $2^{H_{\infty}(S)}\geq 2^{\kappa}$, so Grover’s complexity is $\Omega(2^{\kappa/2})$, providing $\kappa/2$-bit quantum security.
Quantum Collision Attacks: BHT algorithm [7] finds collisions in time $\mathcal{O}(2^{m/3})$. Setting $m\geq 3\kappa$ ensures $2^{m/3}\geq 2^{\kappa}$.
Quantum Memory Attacks: Adversaries storing quantum states require [21, 27]:
$\gamma\geq\kappa+\log Q+\log(1/\epsilon).$ (33)
Our bounded quantum memory assumption $Q\leq 2^{\kappa/2}$ ensures feasibility with $\gamma=\kappa+\kappa/2=1.5\kappa$.
Quantum Rewinding Attacks: Addressed by QUC simulator’s extraction technique [6, 31].
Parameterization for 128-bit Security:
$n$ $=5,\quad m=384,\quad\delta=10,\quad\epsilon=2^{-40}$
$\gamma$ $=\left\lceil\frac{128+40+5\times 10+4\times 384}{5}\right\rceil$
$=\left\lceil\frac{128+40+50+1536}{5}\right\rceil=\lceil 350.8\rceil=351$
$H_{\infty}(S)$ $\geq 5\times(351-10)-4\times 384$
$=5\times 341-1536=1705-1536=169\geq 168.$
The protocol’s resilience to quantum attacks stems from its layered defense strategy:
  • Structural defense: The XOR operation’s linear algebra properties prevent quantum speedup exploitation
  • Entropic defense: Min-entropy bounds ensure exponential search spaces
  • Cryptographic defense: QROM-based commitments resist quantum collision attacks
Theorem 9 (Comprehensive Quantum Security).
For any QPT adversary $\mathscr{A}$ with quantum memory $Q=2^{q}$ qubits, the distinguishing advantage satisfies:
$\left|\Pr[\mathscr{A}(K)=1]-\Pr[\mathscr{A}(U)=1]\right|$
$\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\frac{n^{2}}{2^{\lambda}}+2^{q-(n\gamma-(n-1)m)}.$ (34)
Theorem 9 provides a unified security bound incorporating all quantum attack vectors. The $2^{-\kappa}$ term represents the ideal key randomness, $\frac{q^{2}}{2^{m}}$ bounds commitment collisions, $\frac{n^{2}}{2^{\lambda}}$ covers authentication forgeries, and $2^{q-(n\gamma-(n-1)m)}$ quantifies quantum memory attacks. Our parameterization ensures all terms are $\leq 2^{-128}$ for 128-bit security.
5-C1 Quantum Collision Resistance Analysis
The Brassard-Høyer-Tapp (BHT) attack [7] achieves $O(2^{m/3})$ complexity by:
[…]
$\text{Measure collision with prob. }p\geq c\cdot 2^{-m/3}.$
Our parameterization $m=3\kappa$ ensures:
$\text{Expected queries}=\sqrt{\frac{\pi}{4p}}\geq\sqrt{\frac{\pi}{4c}}\cdot 2^{\kappa}\gg 2^{\kappa}.$ (35)
The commitment structure $c_{i}=\mathcal{H}(s_{i}\parallel\hat{H}_{i})$ forces domain separation, preventing Wagner’s generalized birthday attacks in quantum settings.
6 Entropy Requirements and Parameter Trade-offs
6-A Theoretical Parameter Optimization
Optimize parameters for security and efficiency ($\kappa=128$, $\delta=10$, $m=384$, $\epsilon=2^{-40}$):
Table 2 incorporates 16-bit security margins to mitigate estimation uncertainties and unforeseen attacks. The margin $H_{\infty}(S)-\kappa$ absorbs potential reductions from quantum side information $S_{0}(\rho_{E})$. For $n=4$, $\gamma$ increases to 340 ensuring $H_{\infty}(S)\geq 152>\kappa+24$, satisfying:
$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)+\zeta$ (36)
where $\zeta=24$ represents the operational security buffer. This conservative parameterization accounts for possible deviations in entropy estimation and quantum memory effects.
TABLE 2: Parameterization for $\kappa=128$-bit Quantum Security
$n$ $m$ $\gamma$ $H_{\infty}(S)$ Comm. Cost (KB) Security Margin
3 384 315 135 0.42 7
4 384 340 168 0.84 40
5 384 351 169 1.41 41
6 384 352 172 2.25 44
7 384 359 179 3.15 51
Note: Communication cost computed as $\frac{n(n-1)m}{8\times 1024}$ KB
Design Guidelines:
  • Small $n$: Higher $\gamma$ required, but lower communication
  • Large $n$: Lower $\gamma$ possible, but quadratic communication overhead
  • Balanced: $n=5$ provides optimal tradeoff for 128-bit security
  • Security margin: $H_{\infty}(S)-\kappa$ provides buffer against unforeseen attacks
6-A1 Quantum-Secure Entropy Estimation
Conventional min-entropy estimators exhibit bias under quantum sampling. Adopting quantum-proof techniques [33], we bound estimation error:
Theorem 10 (Quantum Min-Entropy Sampling).
For $\epsilon>0$ and samples $\mathcal{S}=\{x_{1},\dots,x_{N}\}$ from source $X$:
$\Pr_{x_{i}\leftarrow X}\left[\hat{H}_{\infty}(X)\geq S_{\infty}(X|E)_{\rho}-\Delta\right]\geq 1-\epsilon$ (37)
with $\Delta=\log(1/\epsilon)+\frac{1}{2}\log|\mathcal{X}|-H_{2}(X)$. When $N\geq\frac{2}{\epsilon^{2}}\log(2|\mathcal{X}|)$, $\Delta\leq 2\delta$ for $\delta$ in Definition 3.2.
The parameter $\delta$ in our protocol absorbs $\Delta$, ensuring $\gamma-\delta$ remains a reliable lower bound even against quantum-advantaged estimation attacks. This adaptivity is crucial for maintaining composable security under quantum side information.
6-B Quantum Advantage Mitigation
To counter quantum speedups:
  • Grover mitigation: Set $\kappa^{\prime}=2\kappa$ for 128-bit quantum security
  • BHT mitigation: Set $m\geq 3\kappa$
  • Quantum memory: Set $\gamma\geq\kappa+\log Q$
  • Error margin: Include $\delta$ buffer for entropy estimation
  • Composability: Use QUC framework for modular security [6]
6-B1 Quantum Effects on Entropy Estimation
Conventional min-entropy estimators exhibit vulnerabilities under quantum sampling:
$\sup_{\rho}\left|\hat{H}_{\infty}^{\text{class}}(X)-S_{\infty}(X|E)_{\rho}\right|\leq\delta+\log(1/\epsilon)$ (38)
Our solution employs quantum-proof estimators via:
  1. 1.
    Quantum-secure randomness extractors: $\text{Ext}:\{0,1\}^{m}\times\{0,1\}^{d}\rightarrow\{0,1\}^{\kappa}$
  2. 2.
    Two-universal hashing with quantum side information
  3. 3.
    Min-entropy sampling from quantum sources [15]:
    $\Pr_{x^{n}\leftarrow X^{n}}\left[\hat{H}_{\infty}(X)\geq H_{\infty}(X)-\Delta\right]\geq 1-\epsilon$ (39)
    where $\Delta=O(\sqrt{n}/|\mathcal{X}|)$.
The $\delta$ parameter absorbs quantum sampling errors, maintaining $\gamma-\delta$ as effective min-entropy.
7 Comparative Analysis and Protocol Extensions
7-A Comparative Analysis with Post-Quantum Alternatives
The quantum security landscape features diverse approaches with fundamentally different security foundations. Our protocol’s distinctive information-theoretic security provides unique advantages compared to computational post-quantum solutions. As shown in Table 3, while lattice-based schemes like CRYSTALS-Kyber offer efficient $\mathcal{O}(n\kappa)$ communication, their security relies on the unproven hardness of module-LWE problems, which remain vulnerable to unforeseen quantum algorithmic breakthroughs. Similarly, isogeny-based schemes like SIKE provide compact key sizes but have suffered devastating cryptanalytic attacks in recent years, demonstrating the fragility of dependency on specific mathematical assumptions.
TABLE 3: Post-Quantum Key Agreement Protocol Comparison
Protocol Security Basis Sec. Type Comm. Comp. Hardware Trust Hardness
Our Protocol Rényi entropy Info-theoretic $\mathcal{O}(n^{2}m)$ $\mathcal{O}(n^{2})$ Classical Honest majority None
CRYSTALS-Kyber [8] Module-LWE Computational $\mathcal{O}(n\kappa)$ $\mathcal{O}(n\kappa^{2})$ Classical PKI MLWE
SPHINCS+ [28] Hash functions QROM $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa)$ Classical None QROM
SIKE [11] Isogenies Computational $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa^{4})$ Classical PKI SSI
QKD [25] Quantum physics Info-theoretic $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa)$ Quantum Auth. chan. None
Sec. Type: Security Type; Comm.: Communication Complexity; Comp.: Computational Complexity;
Hardware: Hardware Requirements; Trust: Trust Assumptions; Hardness: Hardness Assumptions.
Auth. chan.: Authenticated channels; PKI: Public Key Infrastructure.
Quantum key distribution (QKD) shares our information-theoretic security properties but requires specialized quantum communication hardware and authenticated classical channels. Our protocol achieves comparable security using classical channels only, making it deployable in existing network infrastructure. Crucially, our approach provides built-in fault tolerance against malicious participants through the t1t-1 threshold security of secret sharing, a feature absent in both computational PQC and QKD systems.
The most significant advantage is our protocol’s provable min-entropy guarantee $H_{\infty}(K)\geq\kappa$, which ensures security even against future quantum algorithmic advances. This quantifiable security metric provides long-term assurance unavailable in computational approaches. For $\kappa=128$-bit security with $n=5$ parties, we achieve $H_{\infty}(K)\geq 169$ bits, providing a security buffer against unforeseen attacks.
7-B Protocol Extensions to Secure Applications
The core protocol naturally extends to several high-impact applications through novel cryptographic frameworks:
1. Secure Multiparty Computation for Linear Functions: The protocol directly supports privacy-preserving computation of linear functions $f(s_{1},\dots,s_{n})=\sum c_{i}s_{i}$ without revealing individual inputs. This enables:
  • Federated learning: Secure aggregation of model updates while preserving data privacy
  • Financial auditing: Cross-institutional fraud detection with confidential inputs
  • Supply chain optimization: Collaborative logistics planning with proprietary data protection
The derived key $K_{f}=\mathcal{H}(f(s_{1},\dots,s_{n}))$ maintains the min-entropy guarantee $H_{\infty}(K_{f})\geq\kappa$ when coefficients $c_{i}$ are properly constrained.
2. Quantum-Hybrid Security Architecture: Integrating with quantum key distribution creates a defense-in-depth architecture:
$\text{Hybrid Key}=\mathcal{H}(K_{\text{QKD}}\oplus K_{\text{Entropy}})$ (40)
This combines the active security of QKD with the fault tolerance of our entropy protocol, creating a quantum-resistant solution suitable for critical infrastructure. The hybrid approach provides:
  • Enhanced active security through QKD’s authentication mechanisms
  • Fault tolerance against malicious participants via secret sharing
  • Defense in depth where compromise of one system doesn’t break overall security
3. Post-Quantum Authentication Framework: We develop an entropy-based message authentication code (MAC):
$\text{MAC}(k,m,s)=\mathcal{H}(k\oplus\mathcal{H}(m\|s))$ (41)
leveraging the same entropy sources used in key agreement. This provides quantum-resistant authentication with security bound $\Pr[\text{forge}]\leq\frac{q^{2}}{2^{\gamma}}+\frac{(q+1)^{2}}{2^{\lambda}}$, integrating seamlessly with our key establishment protocol.
7-C Future Research Directions
Four high-impact research directions emerge from this work:
1) Lightweight Implementations for IoT: Developing optimized implementations for resource-constrained devices presents significant challenges. The quadratic communication complexity $\mathcal{O}(n^{2}m)$ becomes problematic for large $n$, requiring compression techniques for shares and commitments. Research should explore:
  • Hierarchical secret sharing to reduce communication
  • Efficient entropy estimation with limited samples
  • Hardware acceleration for polynomial operations
2) Blockchain Integration for Randomness Beacons: The protocol can power decentralized randomness generation for blockchain consensus:
$K_{\text{epoch}}=\mathcal{H}\left(\bigoplus_{i=1}^{n}\mathcal{H}(s_{i}\|\text{block}_{h-1})\right)$ (42)
providing verifiable randomness with min-entropy bound $H_{\infty}(K_{\text{epoch}}|\mathcal{B})\geq\gamma-\log(n|\mathcal{B}|)$. This enables:
  • Bias-resistant consensus protocols
  • Fair NFT minting and airdrops
  • Transparent on-chain gambling
3) Post-Quantum Cryptography Standardization: Our work provides foundations for standardization efforts in information-theoretic PQC. Key initiatives include:
  • Formal security proofs using quantum proof assistants
  • Parameterization guidelines for different security levels
  • Implementation testing frameworks Side-channel resistance certification
4) Entropy Source Diversity and Robustness: Future work must address practical challenges in entropy source management:
  • Security against adversarial entropy sources
  • Cross-source min-entropy estimation techniques
  • Quantum-resistant entropy pooling methods
  • Continuous entropy validation during operation
These extensions and research directions significantly broaden the protocol’s applicability while maintaining its core security properties, positioning it as a foundational element for long-term quantum-resistant cryptography.
8 Conclusion
We have presented a theoretical framework for post-quantum key agreement based on Rényi entropy. The enhanced protocol addresses vulnerabilities in prior constructions through innovations: (1) a confidentiality-preserving verification mechanism using distributed polynomial commitments, (2) provably non-negative min-entropy bounds for XOR composition, and (3) composable security proofs in the quantum universal composability model.
The protocol’s security rests on three mathematical pillars derived from quantum information theory:
Entropy Preservation: $H_{\infty}(S)$ $\geq n\gamma-(n-1)m$ (43)
Verification Security: $\Pr[\text{bypass}]$ $\leq 2^{-\kappa}$ (44)
Composable Security: $\delta_{\text{QUC}}$ $\leq\operatorname{negl}(\kappa)$ (45)
Equations (43)-(45) form an integrated security framework that resists quantum attacks through fundamental information-theoretic principles rather than computational assumptions. For 128-bit security, our parameter optimization yields:
$\boxed{n=5,\quad m=384,\quad\gamma=351\quad\Rightarrow\quad H_{\infty}(K)\geq 169}$
with communication overhead $\mathcal{O}(n^{2}m)=1.41$ KB - a practical cost for long-term security.
Theoretical analysis demonstrates information-theoretic security against passive quantum adversaries and active security with authenticated channels.
Key advantages include:
  • Information-theoretic security without reliance on hardness assumptions
  • Resistance to quantum algorithm breakthroughs
  • Fault tolerance via secret sharing
  • Extensibility to multiparty computation and hybrid systems
Future work includes developing lightweight variants for IoT applications, formal verification using quantum proof assistants, and standardization efforts. By leveraging information-theoretic principles, this work establishes a paradigm for long-term cryptographic security in the quantum era.
Biography Section
Ruopengyu Xu received his B.S degree in network engineering from Software Engineering Institute Gungzhou in June 2025. He is currently an independent scientific researcher. Recently, he has published a paper in https://arxiv.org/abs/2506.07640. At the same time, he is also preparing his postgraduate plan. His research interests are in Quantum cryptography, Information Security, Number Theory and Public key cryptosystem so on.
Chenglian Liu received his B.S. degree in information management from National Union University in 1992, and his M.Sc. degree in National Defense from National Defense University in 2004. He was a research fellow in the Department of Automation Engineering at Technological Educational Institution of Sterea Hellas from 2017 to 2019, and a visiting scholar at Universidad del Pacifico from 2020 to 2022. He is currently an associate professor at the School of Electrical and Computer Engineering, Nanfang University, Guangzhou, China. His main research focuses on Information Security, Cryptanalysis, Fuzzy Sets, and Goldbach’s Conjecture, among other topics.

### v2

@@
-Enhanced Rényi Entropy-Based Post-Quantum Key Agreement with Provable Security and Information-Theoretic Guarantees
+Enhanced Renyi Entropy-Based Post-Quantum Key Agreement with Provable Security and Information-Theoretic Guarantees
 Abstract
-This paper presents an enhanced post-quantum key agreement protocol based on Rényi entropy, addressing vulnerabilities in the original construction while preserving information-theoretic security properties.
+This paper presents an enhanced post-quantum key agreement protocol based on Renyi entropy, addressing vulnerabilities in the original construction while preserving information-theoretic security properties.
 We develop a theoretical framework leveraging entropy-preserving operations and secret-shared verification to achieve provable security against quantum adversaries.
@@
 Extensions to secure multiparty computation and quantum network applications are established, providing a foundation for long-term cryptographic security.
-All security claims are derived from mathematical proofs; this theoretical work presents no experimental validation.
-Index Terms:
-Rényi entropy, post-quantum cryptography, key agreement, information-theoretic security, quantum security, entropy preservation, secret sharing, quantum random oracle model.
+keywords:
+Renyi entropy , post-quantum cryptography , key agreement , information-theoretic security , quantum security , entropy preservation , secret sharing , quantum random oracle model
+journal: Theoretical Computer Science
+[…]
+organization=Independent Researcher,city=,country=
+[…]
+organization=School of Electrical and Computer Engineering, Nanfang College Guangzhou,addressline=, city=Guangzhou, postcode=510970, country=China
 1 Introduction
-The anticipated development of large-scale quantum computing represents a significant challenge to contemporary cryptographic infrastructure [31, 20, 14, 1].
+The anticipated development of large-scale quantum computing represents a significant challenge to contemporary cryptographic infrastructure [1, 20, 14].
 Shor’s polynomial-time quantum algorithm for integer factorization and discrete logarithms [1] compromises the security of widely-deployed asymmetric cryptosystems including RSA, ECC, and Diffie-Hellman key exchange.
@@
 This quantum threat landscape has catalyzed post-quantum cryptography (PQC) development, with NIST standardizing lattice-based, code-based, and multivariate schemes [20, 9].
-However, these approaches rely on computational hardness assumptions that may be vulnerable to algorithmic advances [26, 28, 10].
-1-A Research Context and Motivation
+However, these approaches rely on computational hardness assumptions that may be vulnerable to algorithmic advances [10, 28].
+1.1 Research Context and Motivation
 Information-theoretic cryptography offers an alternative paradigm with unconditional security guarantees that persist against quantum adversaries possessing unbounded computational resources [15, 21].
@@
 Classical information-theoretic solutions based on secret sharing [4] and physical unclonable functions exist but typically require pre-shared keys or lack quantum resistance [17].
-The original Rényi entropy-based key agreement protocol [32] demonstrated theoretical promise but contained vulnerabilities: (1) input exposure during broadcast enabling key compromise, (2) potentially negative entropy bounds violating security requirements, and (3) insufficient protection against quantum-specific attacks.
+The original Renyi entropy-based key agreement protocol [32] demonstrated theoretical promise but contained vulnerabilities: (1) input exposure during broadcast enabling key compromise, (2) potentially negative entropy bounds violating security requirements, and (3) insufficient protection against quantum-specific attacks.
 This work addresses these limitations through three key innovations:
@@
 The theoretical foundation of our approach rests on three interconnected pillars of quantum information theory: (1) Quantum entropy preservation - ensuring min-entropy bounds hold against quantum adversaries with side information; (2) Distributed verification - enabling secure consistency checks without exposing sensitive inputs; (3) Composable security - providing security guarantees under quantum composition.
-Theorem 1 (Quantum Entropy Preservation Bound).
+Theorem 1.1 (Quantum Entropy Preservation Bound).
 For $n$ independent entropy sources $s_{i}$ with $H_{\infty}(s_{i})\geq\gamma$ and quantum adversary holding quantum side information $\rho_{E}$, the combined secret $S=\bigoplus_{i=1}^{n}s_{i}$ satisfies:
-$S_{\infty}(S|\rho_{E})\geq n\gamma-(n-1)m-S_{0}(\rho_{E})$ (1)
+$S_{\infty}(S|\rho_{E})\geq n\gamma-(n-1)m-S_{0}(\rho_{E})$
 where $S_{0}(\rho_{E})$ is the quantum max-entropy of $\rho_{E}$.
@@
 The $S_{0}(E)$ term quantifies the security degradation from quantum memory attacks, which we mitigate through parameter optimization.
-1-B Key Contributions
+1.2 Key Contributions
 This paper makes seven significant advances in post-quantum cryptography:
-1.
+  1. 1.
     Confidentiality-Preserving Verification: A secret-shared verification mechanism using distributed polynomial commitments that prevents input exposure while enabling secure entropy verification, addressing a vulnerability in prior constructions [32].
-2. Optimal Entropy Bounds: Mathematically rigorous min-entropy preservation theorems with guaranteed non-negative bounds and parameterization strategies for quantum security.
-3.
+  2. 2.
+    Optimal Entropy Bounds: Mathematically rigorous min-entropy preservation theorems with guaranteed non-negative bounds and parameterization strategies for quantum security.
+  3. 3.
     Quantum Universal Composability: Composable security proof for information-theoretic key agreement in the quantum universal composability (QUC) model [5, 6], demonstrating secure realization of ideal key functionality.
-4.
+  4. 4.
     Quantum Attack Resistance: Formal analysis against known quantum attack vectors including Grover-accelerated brute force, quantum collision search [7], and quantum memory attacks [21], with quantifiable security bounds.
-5. Entropy Amplification Framework: Generalized framework for multi-party entropy amplification using R’enyi entropy measures [3], providing exponential security scaling against quantum adversaries.
-6. Hybrid Security Extension: Integration with quantum key distribution (QKD) [25] that enhances security against active adversaries while preserving information-theoretic guarantees.
-7. Secure Computation Extension: Secure extension to general secure multiparty computation [24, 17] for linear functions, enabling privacy-preserving applications.
+  5. 5.
+    Entropy Amplification Framework: Generalized framework for multi-party entropy amplification using R’enyi entropy measures [3], providing exponential security scaling against quantum adversaries.
+  6. 6.
+    Hybrid Security Extension: Integration with quantum key distribution (QKD) [25] that enhances security against active adversaries while preserving information-theoretic guarantees.
+  7. 7.
+    Secure Computation Extension: Secure extension to general secure multiparty computation [17, 24] for linear functions, enabling privacy-preserving applications.
 These contributions establish a paradigm for quantum-resistant cryptography based on information-theoretic principles, with applications in secure multiparty computation and quantum networks.
-1-C Quantum Entropic Framework
+1.3 Quantum Entropic Framework
 The foundation of our approach rests upon the rigorous quantification of uncertainty in quantum systems.
-Unlike classical entropy measures, quantum Rényi entropy (9) captures the fundamental limits of information extraction under quantum mechanical constraints.
+Unlike classical entropy measures, quantum Renyi entropy (4) captures the fundamental limits of information extraction under quantum mechanical constraints.
 This becomes crucial when adversaries possess quantum memory capable of storing superposition states for delayed measurement [21].
 Our framework explicitly addresses this quantum advantage by establishing composable entropy bounds through:
+  • 1.
     Quantum-proof min-entropy extraction from non-uniform sources
+  • 2.
     Entropic uncertainty relations under quantum side information
+  • 3.
     Tight bounds on quantum guessing probability via $S_{\infty}(\rho)$
@@
 2 Background and Theoretical Foundations
-2-A Quantum Computing Threat Model
+2.1 Quantum Computing Threat Model
 Quantum computation utilizes state superposition and quantum correlation effects to achieve computational advantages for specific problems [14]. We formalize the quantum adversary model:
-Definition 1 (Quantum Polynomial-Time Adversary).
-A quantum adversary $\mathscr{A}$ is a polynomial-time quantum algorithm with:
+Definition 2.1 (Quantum Polynomial-Time Adversary).
+A quantum adversary $\mathcal{A}$ is a polynomial-time quantum algorithm with:
+  • 1.
     Quantum random oracle access to hash functions
+  • 2.
     Quantum memory bounded by $Q=2^{\mathcal{O}(\kappa)}$ qubits
+  • 3.
     Capability to corrupt up to t1t-1 parties
+  • 4.
     Ability to perform superposition queries to oracles
+  • 5.
     Adaptive measurement strategies [21, 31]
 Threat Analysis:
+  • 1.
     Shor’s Algorithm: Factors integers in $\mathcal{O}((\log N)^{3})$ time [1], compromising RSA, ECC, and Diffie-Hellman
+  • 2.
     Grover’s Algorithm: Solves unstructured search in $\mathcal{O}(\sqrt{N})$ time [2], reducing symmetric key security
-  • Brassard-Høyer-Tapp (BHT): Quantum collision finding in $\mathcal{O}(2^{m/3})$ time [7]
+  • 3.
+    Brassard-Hoyer-Tapp (BHT): Quantum collision finding in $\mathcal{O}(2^{m/3})$ time [7]
+  • 4.
     Quantum Memory Attacks: Exploit coherent state persistence in quantum storage [21, 30]
+  • 5.
     Quantum Rewinding: Subvert classical proof systems [31]
 Research Significance: Our protocol design counters these threats through entropy amplification and information-theoretic security [30, 21], ensuring security against quantum adversaries.
-2-B Rényi Entropy Framework
-Rényi entropy [3] provides a parametric family of entropy measures essential for cryptographic security analysis.
+2.2 Renyi Entropy Framework
+Renyi entropy [3] provides a parametric family of entropy measures essential for cryptographic security analysis.
 Its operational significance in quantum cryptography stems from connections to guessing probabilities [18]:
-Definition 2 (Rényi Entropy).
-For discrete random variable $X\sim P_{X}$ over $\mathcal{X}$, Rényi entropy of order $\alpha$ is:
-$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\left(\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}\right)$ (2)
+Definition 2.2 (Renyi Entropy).
+For discrete random variable $X\sim P_{X}$ over $\mathcal{X}$, Renyi entropy of order $\alpha$ is:
+$H_{\alpha}(X)=\frac{1}{1-\alpha}\log\left(\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}\right)$
 for $\alpha>0$, $\alpha\neq 1$.
 Cryptographically significant special cases:
-$\lim_{\alpha\to 1}H_{\alpha}(X)=H(X)=-\sum_{x}P_{X}(x)\log P_{X}(x)$ (Shannon) (3)
-$H_{2}(X)=-\log\left(\sum_{x}P_{X}(x)^{2}\right)$ (Collision) (4)
-$\lim_{\alpha\to\infty}H_{\alpha}(X)=H_{\infty}(X)=-\log\max_{x}P_{X}(x)$ (Min-entropy) (5)
+$\lim_{\alpha\to 1}H_{\alpha}(X)$ $=H(X)=-\sum_{x}P_{X}(x)\log P_{X}(x)\quad\text{(Shannon)}$
+$H_{2}(X)$ $=-\log\left(\sum_{x}P_{X}(x)^{2}\right)\quad\text{(Collision)}$
+$\lim_{\alpha\rightarrow\infty}H_{\alpha}(X)$ $=H_{\infty}(X)=-\log\max_{x}P_{X}(x)\quad\text{(Min-entropy)}$
 Operational Significance [15, 18, 19]:
+  • 1.
     Min-entropy: $H_{\infty}(X)=-\log P_{\text{guess}}(X)$ where $P_{\text{guess}}$ is optimal guessing probability
+  • 2.
     Collision entropy: $H_{2}(X)=-\log P_{\text{coll}}(X)$ where $P_{\text{coll}}$ is collision probability
+  • 3.
     Quantum advantage: Min-entropy bounds quantum guessing probability under side information
 Fundamental properties enabling our security proofs:
-Lemma 1 (Entropy Transformation).
+Lemma 2.3 (Entropy Transformation).
 For independent $X,Y$ and deterministic function $f$:
@@
 Follows from probability bound $\max_{z}\Pr[f(X)=z]\geq\frac{\max_{x}\Pr[X=x]}{|\text{range}(f)|}$ and convolution properties. Detailed proof in [13]. ∎
-2-C Quantum Information Theoretic Foundations
-Definition 3 (Classical vs. Quantum Entropy Notation).
+2.3 Quantum Information Theoretic Foundations
+Definition 2.4 (Classical vs. Quantum Entropy Notation).
 Throughout this paper:
-  • $H_{\alpha}(X)$ denotes classical Rényi entropy for random variable $X$
-  • $S_{\alpha}(\rho)$ denotes quantum Rényi entropy for density operator $\rho$
+  • 1.
+    $H_{\alpha}(X)$ denotes classical Renyi entropy for random variable $X$
+  • 2.
+    $S_{\alpha}(\rho)$ denotes quantum Renyi entropy for density operator $\rho$
+  • 3.
     $H_{\infty}(X)$ and $S_{\infty}(\rho)$ specifically denote min-entropy
 For quantum systems, security analysis requires quantum entropy measures [19]:
-Definition 4 (Quantum Rényi Entropy).
-For density operator $\rho$, quantum Rényi entropy is:
-$S_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\operatorname{tr}(\rho^{\alpha})$ (9)
+Definition 2.5 (Quantum Renyi Entropy).
+For density operator $\rho$, quantum Renyi entropy is:
+$S_{\alpha}(\rho)=\frac{1}{1-\alpha}\log\textnormal{tr}(\rho^{\alpha})$ (4)
 Quantum Security Bounds [15, 30, 21, 23]:
+  • 1.
     Quantum min-entropy: $S_{\infty}(\rho)=-\log\lambda_{\max}(\rho)$ bounds state guessing probability
-  • Quantum collision entropy: $S_{2}(\rho)=-\log\operatorname{tr}(\rho^{2})$ quantifies state distinguishability
+  • 2.
+    Quantum collision entropy: $S_{2}(\rho)=-\log\textnormal{tr}(\rho^{2})$ quantifies state distinguishability
+  • 3.
     Data processing inequality: $S_{\alpha}(\mathscr{E}(\rho))\geq S_{\alpha}(\rho)$ for quantum channels $\mathscr{E}$
+  • 4.
     Entropy accumulation: $S_{\infty}(\rho_{X^{n}E})\geq\sum_{i=1}^{n}S_{\infty}(\rho_{X_{i}E})-c\sqrt{n}$ [23]
 Research Significance: Our security proofs leverage these quantum entropy measures to establish composable security against quantum adversaries.
-2-D Distributed Verification via Secret Sharing
+2.4 Distributed Verification via Secret Sharing
 Shamir’s secret sharing [4] provides the foundation for our confidentiality-preserving verification:
-Theorem 2 (Shamir’s Secret Sharing).
+Theorem 2.6 (Shamir’s Secret Sharing).
 For secret $s\in\mathbb{F}_{q}$, threshold $t$, and $n$ parties, choose random polynomial:
@@
 Distribute shares $s_{i}=f(i)$. Then:
+  • 1.
     Any $t$ shares reconstruct $s$ via Lagrange interpolation
+  • 2.
     Any t1t-1 shares reveal zero information about $s$
 We enhance this scheme with verifiable features [24]:
+  • 1.
     Polynomial commitments: Binding cryptographic commitments to coefficients
+  • 2.
     Distributed verification: Consistency checks without reconstruction
+  • 3.
     Information-theoretic confidentiality: Guaranteed by secret sharing properties
-2-E Symbol Table
-TABLE 1: Mathematical Symbols and Definitions
-Symbol Meaning
-$n$ Number of protocol participants
-$\kappa$ Security parameter (bit strength)
-$m$ Bit-length of secret shares
-$\gamma$ Min-entropy threshold per share
-$t$ Reconstruction threshold ($t=\lfloor n/2\rfloor+1$)
-$\delta$ Entropy estimation accuracy
-$H_{\alpha}(X)$ Classical Rényi entropy of order $\alpha$
-$S_{\alpha}(\rho)$ Quantum Rényi entropy
-$H_{\infty}(X)$ Min-entropy: $\log\max_{x}P_{X}(x)$
-$\hat{H}_{i}$ Estimated min-entropy of $s_{i}$
-$s_{i}$ Secret share generated by party $i$
-$c_{i}$ Cryptographic commitment: $\mathcal{H}(s_{i}\|\hat{H}_{i})$
-$f_{i}(x)$ Sharing polynomial for $s_{i}$
-$sh_{ij}$ Share of $s_{i}$ distributed to $P_{j}$
-$\mathbb{F}_{2^{m}}$ Finite field with $2^{m}$ elements
-$\mathcal{H}$ Quantum random oracle
-$\oplus$ Bitwise XOR operation
-$Q$ Adversary quantum memory bound ($Q=2^{\kappa/2}$ qubits)
-$\mathscr{A}$ Quantum polynomial-time adversary
-$\mathcal{F}_{KA}$ Ideal key agreement functionality
-$K$ Established secret key
-$\lambda$ Commitment security parameter ($\lambda=2\kappa$)
-$\epsilon$ Security error bound ($\epsilon=2^{-40}$)
 3 Enhanced Protocol Design
-3-A System Model and Threat Analysis
-Consider a network of $n$ parties $\mathcal{P}=\{P_{1},\dots,P_{n}\}$ establishing shared secret key $K\in\{0,1\}^{\kappa}$.
+3.1 System Model and Threat Analysis
+Consider a network of $n$ parties $\mathcal{P}=\{P_{1},\ldots,P_{n}\}$ establishing shared secret key $K\in\{0,1\}^{\kappa}$.
 Adversarial Capabilities [6, 21, 31]:
+  • 1.
     Quantum polynomial-time computation
-  • Quantum random oracle access to $\mathcal{H}$
+  • 2.
+    Quantum random oracle access to hash functions
+  • 3.
     Quantum memory bounded by $Q=2^{\kappa/2}$ qubits
+  • 4.
     Adaptive corruption of up to t1t-1 parties
+  • 5.
     Control over communication channels
+  • 6.
     Superposition queries and delayed measurement
+  • 7.
     Quantum rewinding attacks
 Security Assumptions:
+  • 1.
     Authenticated secure channels between parties
+  • 2.
     Reliable entropy sources with $\delta$-accurate min-entropy estimation
+  • 3.
     Quantum-resistant hash function $\mathcal{H}$ modeled as QROM
+  • 4.
     Honest majority ($t=\lfloor n/2\rfloor+1$)
+  • 5.
     Entropy independence across parties
 Design Rationale: This model balances practical quantum threats with theoretical tractability, enabling security proofs while capturing essential quantum capabilities.
-3-B Confidentiality-Preserving Verification Mechanism
+3.2 Confidentiality-Preserving Verification Mechanism
 The original vulnerability stemmed from broadcasting $s_{i}$, allowing adversaries to compute $K=\mathcal{H}({\oplus}_{i}s_{i})$. Our solution integrates secret sharing with cryptographic commitments:
-Definition 5 (Distributed Polynomial Commitment).
+Definition 3.1 (Distributed Polynomial Commitment).
 For secret $s_{i}$, generate random polynomial:
 $f_{i}(x)=s_{i}+\sum_{k=1}^{t-1}a_{k}x^{k}\in\mathbb{F}_{2^{m}}[x]$ (6)
-with $a_{k}\xleftarrow{\$}\{0,1\}^{m}$. The commitment is $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ where $\hat{H}_{i}\approx H_{\infty}(s_{i})$.
+with $a_{k}\overset{\mathdollar}{\leftarrow}\{0,1\}^{m}$. The commitment is $c_{i}=\mathcal{H}(s_{i}||\hat{H}_{i})$ where $\hat{H}_{i}\approx H_{\infty}(s_{i})$.
 Verification Protocol:
@@
   2. 2.
-    $P_{j}$ collects $t$ shares $\{sh_{ik}\}_{k\in S}$
+    $P_{j}$ collects $t$ shares $\{sh_{ik}\}_{k\in\mathcal{S}}$
   3. 3.
@@
   4. 4.
-    Verifies $c_{i}=\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})$
+    Verifies $c_{i}=\mathcal{H}(\tilde{s}_{i}||\hat{H}_{i})$
   5. 5.
@@
 Security Properties [4, 6, 24]:
+  • 1.
     Confidentiality: <t<t colluding parties learn nothing about $s_{i}$
+  • 2.
     Binding: Computational binding under QROM
+  • 3.
     Verifiability: Algebraic structure enables consistency checks
+  • 4.
     Quantum resistance: Security preserved under superposition attacks
@@
 This dual-security property is achieved through the algebraic structure of Shamir’s secret sharing, where verification occurs locally on shares without reconstruction.
-Theorem 3 (Verification Security).
+Theorem 3.2 (Verification Security).
 For any QPT adversary $\mathscr{A}$ attempting to submit invalid share $s^{\prime}\neq s$:
-$\Pr[\text{successful verification}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}$ (12)
+$\Pr[\text{successful verification}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}$
 where $q$ is the number of quantum queries to $\mathcal{H}$.
@@
 By setting $m=3\kappa$ and $\gamma>\kappa$, both terms become negligible for $\kappa\geq 128$.
-3-C Key Agreement Protocol
+3.3 Key Agreement Protocol
 The complete protocol operates in four phases with comprehensive quantum resistance:
-Algorithm 1 Enhanced Rényi Key Agreement Protocol
-0: Security parameter $\kappa$, min-entropy threshold $\gamma$, parties $n$, estimation accuracy $\delta$
-0: Shared secret key $K$
- Set $t=\lceil n/2\rceil+1$ {Honest majority threshold}
-Phase 1: Initialization
-for each party $P_{i}\in\mathcal{P}$ do
-  Sample $s_{i}\leftarrow\{0,1\}^{m}$ with $H_{\infty}(s_{i})\geq\gamma$ {High-entropy secret}
-  Estimate $\hat{H}_{i}$ such that $|H_{\infty}(s_{i})-\hat{H}_{i}|\leq\delta$ {Min-entropy estimation}
-  Compute commitment $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ {QROM-based binding}
-  Broadcast $c_{i}$ to all parties {Public commitment}
-end for
-Phase 2: Share Distribution
-for each $P_{i}\in\mathcal{P}$ do
-  Generate random polynomial $f_{i}(x)=s_{i}+\sum_{k=1}^{t-1}a_{k}x^{k}$, $a_{k}\xleftarrow{\$}\{0,1\}^{m}$
-  for each $P_{j}\in\mathcal{P}\setminus\{P_{i}\}$ do
-   Compute share $sh_{ij}=f_{i}(j)$
-   Securely send $(i,sh_{ij})$ to $P_{j}$ via authenticated channel
-  end for
-end for
-Phase 3: Verification
-for each $P_{j}\in\mathcal{P}$ do
-  for each $P_{i}\in\mathcal{P}\setminus\{P_{j}\}$ do
-   Collect $\geq t$ valid shares $\{sh_{ik}\}_{k\in S}$ for $|S|\geq t$
-   Reconstruct $\tilde{s}_{i}=\sum_{k\in S}sh_{ik}\cdot L_{k}(0)$ {Lagrange interpolation}
-   Verify $\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})=c_{i}$ {Commitment consistency}
-   Verify $H_{\infty}(\tilde{s}_{i})\geq\gamma-\delta$ {Entropy threshold}
-   if any verification fails then
-    Abort protocol and output $\bot$
-   end if
-  end for
-end for
-Phase 4: Key Derivation
- Each $P_{i}$ reveals $s_{i}$ to all parties via authenticated channels {Safe after verification}
- Compute $S=\bigoplus_{i=1}^{n}s_{i}$ {Entropy-preserving combination}
- Compute $K=\mathcal{H}(S)$ {Quantum-secure randomness extraction}
-return  $K$
+Algorithm 1 Enhanced Renyi Key Agreement Protocol
+1:Security parameter $\kappa$, min-entropy threshold $\gamma$, parties $n$, estimation accuracy $\delta$
+2:Shared secret key $K$
+3:Set $t=\lceil n/2\rceil+1$ $\triangleright$ Honest majority threshold
+4:
+5:Phase 1: Initialization
+6:for each party $P_{i}\in\mathcal{P}$ do
+7:  Sample $s_{i}\leftarrow\{0,1\}^{m}$ with $H_{\infty}(s_{i})\geq\gamma$ $\triangleright$ High-entropy secret
+8:  Estimate $\hat{H}_{i}$ such that $|H_{\infty}(s_{i})-\hat{H}_{i}|\leq\delta$ $\triangleright$ Min-entropy estimation
+9:  Compute commitment $c_{i}=\mathcal{H}(s_{i}\|\hat{H}_{i})$ $\triangleright$ QROM-based binding
+10:  Broadcast $c_{i}$ to all parties $\triangleright$ Public commitment
+11:end for
+12:
+13:Phase 2: Share Distribution
+14:for each party $P_{i}\in\mathcal{P}$ do
+15:  Generate random polynomial $f_{i}(x)=s_{i}+\sum_{k=1}^{t-1}a_{k}x^{k}$, $a_{k}\overset{\mathdollar}{\leftarrow}\{0,1\}^{m}$
+16:  for each $P_{j}\in\mathcal{P}\setminus\{P_{i}\}$ do
+17:    Compute share $sh_{ij}=f_{i}(j)$
+18:    Securely send $(i,sh_{ij})$ to $P_{j}$ via authenticated channel
+19:  end for
+20:end for
+21:
+22:Phase 3: Verification
+23:for each party $P_{j}\in\mathcal{P}$ do
+24:  for each $P_{i}\in\mathcal{P}\setminus\{P_{j}\}$ do
+25:    Collect $\geq t$ valid shares $\{sh_{ik}\}_{k\in S}$ for $|S|\geq t$
+26:    Reconstruct $\tilde{s}_{i}=\sum_{k\in S}sh_{ik}\cdot L_{k}(0)$ $\triangleright$ Lagrange interpolation
+27:    Verify $\mathcal{H}(\tilde{s}_{i}\|\hat{H}_{i})=c_{i}$ $\triangleright$ Commitment consistency
+28:    Verify $H_{\infty}(\tilde{s}_{i})\geq\gamma-\delta$ $\triangleright$ Entropy threshold
+29:    if any verification fails then
+30:     Abort protocol and output $\bot$
+31:    end if
+32:  end for
+33:end for
+34:
+35:Phase 4: Key Derivation
+36:Each $P_{i}$ reveals $s_{i}$ to all parties via authenticated channels $\triangleright$ Safe after verification
+37:Compute $S=\bigoplus_{i=1}^{n}s_{i}$ $\triangleright$ Entropy-preserving combination
+38:Compute $K=\mathcal{H}(S)$ $\triangleright$ Quantum-secure randomness extraction
+39:return $K$
 Security Justification for Revealing Phase: After successful verification, revealing $s_{i}$ through authenticated channels is secure because: (1) Verification ensures each $s_{i}$ has min-entropy $\geq\gamma-\delta$, (2) The final key $K$ is derived from XOR of all $s_{i}$ followed by hashing, (3) Adversaries cannot alter $s_{i}$ due to authenticated channels, and (4) Commitment binding prevents submission of different values.
 Protocol Properties:
+  • 1.
     Communication Complexity: $\mathcal{O}(n^{2}m)$ bits
+  • 2.
     Computational Complexity: $\mathcal{O}(n^{2})$ field operations
+  • 3.
     Round Complexity: 3 rounds (broadcast, share exchange, key derivation)
+  • 4.
     Fault Tolerance: Resilient to <t<t malicious parties
-3-D Entropy Verification Theoretical Basis
+3.4 Entropy Verification Theoretical Basis
 Theoretical Challenge: Exact min-entropy computation requires complete knowledge of the underlying probability distribution, which is generally infeasible.
@@
     Utilize statistical estimation techniques:
-    $\hat{H}_{\infty}=-\log\left(\max_{x\in\mathcal{S}}\hat{P}(x)\right)\pm\delta$ (13)
+    $\hat{H}_{\infty}=-\log\left(\max_{x\in\mathcal{S}}\hat{P}(x)\right)\pm\delta$
     for empirical distribution $\hat{P}$ over sample $\mathcal{S}$ of size $N\geq 2^{2\gamma}/\epsilon^{2}$ to achieve $\delta$-accuracy with failure probability $\epsilon$
@@
 The $\delta$ parameter must be conservatively chosen to account for statistical uncertainty.
-3-E Quantum Security Design Principles
+3.5 Quantum Security Design Principles
 The protocol architecture incorporates three fundamental quantum-resistant mechanisms:
   1. 1.
-    Superposition Resistance: Polynomial commitments use injective encoding $\mathcal{H}(s_{i}\parallel\hat{H}_{i})$ to prevent quantum ambiguity attacks.
-The concatenation ensures $\forall s_{i}\neq s_{j},\mathcal{H}(s_{i}\parallel\cdot)\cap\mathcal{H}(s_{j}\parallel\cdot)=\emptyset$ with probability $1-\operatorname{negl}(\lambda)$ under QROM.
+    Superposition Resistance: Polynomial commitments use injective encoding $\mathcal{H}(s_{i}\|\hat{H}_{i})$ to prevent quantum ambiguity attacks.
+The concatenation ensures $\forall s_{i}\neq s_{j},\mathcal{H}(s_{i}\|\cdot)\cap\mathcal{H}(s_{j}\|\cdot)=\emptyset$ with probability $1-\mathrm{negl}(\lambda)$ under QROM.
   2. 2.
-[…]
-    $\Delta(\rho_{SE},\rho_{S}\otimes\rho_{E})\leq 2^{-H_{\min}(S|E)}$ (14)
+    Entanglement Breaking: The final XOR operation $S=\bigoplus s_{i}$ acts as a non-commutative operator relative to quantum adversaries’ observation basis.
+For any quantum state $|\psi\rangle=\sum\alpha_{x,y}|x,y\rangle_{E}$, the operation satisfies:
+    $\Delta(\rho_{SE},\rho_{S}\otimes\rho_{E})\leq 2^{-H_{\min}(S|E)}$
     where $\Delta$ is trace distance.
@@
     Quantum Rewinding Protection: The verification phase forces sequential measurement through:
-    Commit $\rightarrow\text{Share}\rightarrow\text{Verify}$
-    $\Downarrow$
-    No-cloning theorem enforced
+    $\text{Commit }\rightarrow\text{Share }\rightarrow\text{Verify}$
     Adversaries cannot maintain superposition beyond verification due to the measurement requirement in Step 3 of Algorithm 1.
 4 Enhanced Theoretical Analysis
-4-A Entropy Amplification Theorem
+4.1 Entropy Amplification Theorem
 The security of key derivation relies on min-entropy preservation during XOR combination:
-Theorem 4 (Min-Entropy Preservation).
-For independent random variables $X_{1},\dots,X_{n}$ over $\{0,1\}^{m}$ with $H_{\infty}(X_{i})\geq\gamma$, the XOR sum $S=\bigoplus_{i=1}^{n}X_{i}$ satisfies:
-$H_{\infty}(S)\geq\max\left(0,n\gamma-(n-1)m\right)\geq\kappa$ (15)
+Theorem 4.1 (Min-Entropy Preservation).
+For independent random variables $X_{1},\ldots,X_{n}$ over $\{0,1\}^{m}$ with $H_{\infty}(X_{i})\geq\gamma$, the XOR sum $S=\bigoplus_{i=1}^{n}X_{i}$ satisfies:
+$H_{\infty}(S)\geq\max\left(0,n\gamma-(n-1)m\right)\geq\kappa$
 Equality holds when $X_{i}$ are uniform and independent.
@@
 $=\max_{y^{\prime}}\Pr[Y=y^{\prime}]\cdot\sum_{x}\Pr[X=x]$
-$=2^{-H_{\infty}(Y)}$ (16)
+$=2^{-H_{\infty}(Y)}$
 However, a tighter bound derives from convolution properties. For finite field $\mathbb{F}_{2^{m}}$, we have:
-$\max_{z}\Pr[X\oplus Y=z]\leq\min\left(2^{-H_{\infty}(X)},2^{-H_{\infty}(Y)}\right)\cdot|\mathcal{X}|\cdot\alpha$ (17)
+$\max_{z}\Pr[X\oplus Y=z]\leq\min\left(2^{-H_{\infty}(X)},2^{-H_{\infty}(Y)}\right)\cdot|\mathcal{X}|\cdot\alpha$
 where $\alpha=\sup_{x,y}\Pr[Y=y|X=x]$. Under independence $\alpha=2^{-H_{\infty}(Y)}$, yielding:
-$H_{\infty}(X\oplus Y)\geq\max\left(H_{\infty}(X),H_{\infty}(Y)\right)+\log|\mathcal{X}|-\beta$ (18)
+$H_{\infty}(X\oplus Y)\geq\max\left(H_{\infty}(X),H_{\infty}(Y)\right)+\log|\mathcal{X}|-\beta$
 with $\beta=\log(1+\operatorname{corr}(X,Y))$.
@@
 This inequality holds because the XOR operation reduces the maximum probability by at most $\log|\mathcal{X}|$ due to the discrete nature of the alphabet. For any $z$, we have:
-$\Pr[X\oplus Y=z]$
-$\leq\max_{x}\Pr[X=x]\cdot\max_{y}\Pr[Y=y]$
-$\leq 2^{-H_{\infty}(X)}\cdot 2^{-H_{\infty}(Y)}$ (20)
+$\Pr[X\oplus Y=z]$ $\leq\max_{x}\Pr[X=x]\cdot\max_{y}\Pr[Y=y]$
+$\leq 2^{-H_{\infty}(X)}\cdot 2^{-H_{\infty}(Y)}$
 Taking the logarithm base 2, we get:
@@
 $\geq-\log(2^{-H_{\infty}(X)}\cdot 2^{-H_{\infty}(Y)})$
-$=H_{\infty}(X)+H_{\infty}(Y).$ (21)
+$=H_{\infty}(X)+H_{\infty}(Y)$
 However, due to the finite alphabet size, the bound is slightly weaker:
-$H_{\infty}(X\oplus Y)\geq H_{\infty}(X)+H_{\infty}(Y)-\log|\mathcal{X}|$ (22)
+$H_{\infty}(X\oplus Y)\geq H_{\infty}(X)+H_{\infty}(Y)-\log|\mathcal{X}|$
 This completes the proof of the convolution inequality.
@@
 $\geq H_{\infty}(S_{n-1})+H_{\infty}(X_{n})-\log|\mathcal{X}|$
-$\geq\left[(n-1)\gamma-(n-2)\log|\mathcal{X}|\right]+\gamma-\log|\mathcal{X}|$
-$=n\gamma-(n-1)\log|\mathcal{X}|.$ (23)
+$\geq[(n-1)\gamma-(n-2)\log|\mathcal{X}|]+\gamma-\log|\mathcal{X}|$
+$=n\gamma-(n-1)\log|\mathcal{X}|.$
 The $\max(0,\cdot)$ ensures non-negativity when $n\gamma<(n-1)\log|\mathcal{X}|$. ∎
@@
 $S_{\infty}(S|E)\geq n\gamma-(n-1)m-q$
-Lemma 2 (Quantum Chain Rule for Min-Entropy).
+Lemma 4.2 (Quantum Chain Rule for Min-Entropy).
 For quantum state $\rho_{X_{1}\cdots X_{n}E}$ and $S=\bigoplus_{i=1}^{n}X_{i}$:
-$S_{\infty}(S|E)\geq\sum_{i=1}^{n}S_{\infty}(X_{i}|X_{1}\cdots X_{i-1}E)-(n-1)\log|\mathcal{X}|.$ (25)
+$S_{\infty}(S|E)\geq\sum_{i=1}^{n}S_{\infty}(X_{i}|X_{1}\cdots X_{i-1}E)-(n-1)\log|\mathcal{X}|.$
 Proof.
@@
 The inductive step follows from recursive application. Critical observation: conditioning on $X_{1}$ does not decrease entropy when $X_{1}$ is independent of $X_{2}E$. ∎
-Corollary 1.
+Corollary 4.3.
 Under pairwise independence: $S_{\infty}(S|E)\geq\sum S_{\infty}(X_{i}|E)-(n-1)\log|\mathcal{X}|$.
-4-B Security Parameterization Framework
-Theorem 5 (Security Parameterization).
+4.2 Security Parameterization Framework
+Theorem 4.4 (Security Parameterization).
 The protocol achieves $\kappa$-bit quantum security if:
-$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$ (26)
+$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$
 for security parameter $\epsilon$, where $\delta$ accounts for entropy estimation error.
@@
 From Theorem 4 and entropy estimation, $H_{\infty}(S)\geq n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$. By the quantum leftover hash lemma [15], for quantum-secure extractor $\mathcal{H}$:
-$\delta(\rho_{KE},\tau_{K}\otimes\rho_{E})\leq\epsilon$ (27)
+$\delta(\rho_{KE},\tau_{K}\otimes\rho_{E})\leq\epsilon$
 where $\delta$ is trace distance, $\tau_{K}$ uniform key, and $\rho_{E}$ quantum side information. Thus $K$ is $\epsilon$-close to uniform independent of $E$. ∎
@@
 s.t. $\quad n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)$
-$\quad m\geq\max(3\kappa,\gamma).$ (28)
+$\quad m\geq\max(3\kappa,\gamma)$
 $\quad\gamma\geq\gamma_{\min}$
 Optimal solution: Set $m=3\kappa$ for BHT resistance, then $\gamma=\frac{\kappa+\log(1/\epsilon)+n\delta+(n-1)m}{n}$.
-4-C Composable Security Framework
+4.3 Composable Security Framework
 We prove security in the quantum universal composability (QUC) model [5, 6, 31]:
-Theorem 6 (Composable Security).
+Theorem 4.5 (Composable Security).
 The protocol $\Pi$ securely realizes the ideal key agreement functionality $\mathcal{F}_{KA}$ in the ($\mathcal{H},\mathcal{AUTH}$)-hybrid model against QPT adversaries.
@@
 Ideal Functionality $\mathcal{F}_{KA}$:
-  • Upon receiving $(\text{init})$ from all honest parties, output $K\xleftarrow{\$}\{0,1\}^{\kappa}$ to all parties
+  • 1.
+    Upon receiving (init) from all honest parties, output $K\overset{\mathdollar}{\leftarrow}\{0,1\}^{\kappa}$ to all parties
+  • 2.
     Adversary learns nothing beyond protocol messages
@@
   1. 1.
-    Commit phase: Generate random $c_{i}^{\prime}\xleftarrow{\$}\{0,1\}^{\lambda}$ without knowing $s_{i}$
+    Commit phase: Generate random $c^{\prime}_{i}\overset{\mathdollar}{\leftarrow}\{0,1\}^{\lambda}$ without knowing $s_{i}$
   2. 2.
-    Share phase: Simulate shares using random polynomials consistent with $c_{i}^{\prime}$
+    Share phase: Simulate shares using random polynomials consistent with $c^{\prime}_{i}$
   3. 3.
-    Verification phase: Program random oracle $\mathcal{H}$ to satisfy $\mathcal{H}(s_{i}^{\prime}\|\hat{H}_{i}^{\prime})=c_{i}^{\prime}$ for extracted $s_{i}^{\prime}$
+    Verification phase: Program random oracle $\mathcal{H}$ to satisfy $\mathcal{H}(s^{\prime}_{i}||\bar{H}^{\prime}_{i})=c^{\prime}_{i}$ for extracted $s^{\prime}_{i}$
   4. 4.
@@
 Indistinguishability: For any environment $\mathcal{Z}$, the distinguishing advantage is bounded by:
-$\left|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{KA},\mathcal{S},\mathcal{Z}}=1]\right|\leq\mathsf{negl}(\kappa)$ (29)
+$|\Pr[\textsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\textsf{IDEAL}_{\mathcal{F}_{KA},\mathcal{S},\mathcal{Z}}=1]|\leq\textsf{negl}(\kappa)$
 This follows from:
+  • 1.
     Indistinguishability of commitments under QROM
+  • 2.
     Information-theoretic secrecy of shares (<t<t parties)
+  • 3.
     Programmability of random oracle [6]
+  • 4.
     Entropy preservation ensuring key uniformity
+  • 5.
     Quantum rewinding security [31]
 5 Security Analysis
-5-A Passive Security Against Quantum Eavesdroppers
-Theorem 7 (Passive Security).
-Under QROM, for any QPT passive adversary $\mathscr{A}$, the distinguishing advantage satisfies:
-$\left|\Pr[\mathscr{A}(K)=1]-\Pr[\mathscr{A}(U)=1]\right|\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\mathsf{negl}(\lambda)$ (30)
+5.1 Passive Security Against Quantum Eavesdroppers
+Theorem 5.1 (Passive Security).
+Under QROM, for any QPT passive adversary $\mathcal{A}$, the distinguishing advantage satisfies:
+$\left|\Pr[\mathcal{A}(K)=1]-\Pr[\mathcal{A}(U)=1]\right|\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\mathsf{negl}(\lambda)$
 where $U$ is uniform random, $q$ is number of quantum queries.
@@
 Adversary advantage decomposes as:
-$\text{Adv}(\mathscr{A})$ $\leq\delta(\text{real protocol},\text{ideal protocol})$
-$+\delta(\text{ideal protocol},\text{uniform}).$ (31)
+$\mathsf{Adv}(\mathcal{A})\leq\delta(\text{real protocol},\text{ideal protocol})+\delta(\text{ideal protocol},\text{uniform}).$
 The first term is negligible by composability (Theorem 3).
@@
 ∎
-5-B Active Adversary Resistance
-Theorem 8 (Active Security).
-With authenticated channels, for any QPT active adversary $\mathscr{A}$:
-$\Pr[\mathscr{A}\text{ wins}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}+\frac{n^{2}}{2^{\lambda}}+\mathsf{negl}(\kappa)$ (32)
+5.2 Active Adversary Resistance
+Theorem 5.2 (Active Security).
+With authenticated channels, for any QPT active adversary $\mathcal{A}$:
+$\Pr[\mathcal{A}\text{ wins}]\leq\frac{q^{2}}{2^{m}}+\frac{1}{2^{\gamma-\delta}}+\frac{n^{2}}{2^{\lambda}}+\mathsf{negl}(\kappa)$
 where win conditions include: (1) forcing acceptance of invalid shares, (2) distinguishing $K$ from random, or (3) causing honest parties to output different keys.
@@
     Share manipulation: Alter $\geq t$ shares without detection, prevented by binding property
-Union bound gives the result. The $\frac{n^{2}}{2^{\lambda}}$ term accounts for commitment forgeries across all parties. ∎
-5-C Quantum Attack Resilience Analysis
+Union bound gives the result. The $\frac{n^{2}}{2^{\kappa}}$ term accounts for commitment forgeries across all parties. ∎
+5.3 Quantum Attack Resilience Analysis
 Grover’s Algorithm Resistance: The search space for $S=\oplus s_{i}$ has size $2^{H_{\infty}(S)}\geq 2^{\kappa}$, so Grover’s complexity is $\Omega(2^{\kappa/2})$, providing $\kappa/2$-bit quantum security.
@@
 Quantum Memory Attacks: Adversaries storing quantum states require [21, 27]:
-$\gamma\geq\kappa+\log Q+\log(1/\epsilon).$ (33)
+$\gamma\geq\kappa+\log Q+\log(1/\epsilon).$
 Our bounded quantum memory assumption $Q\leq 2^{\kappa/2}$ ensures feasibility with $\gamma=\kappa+\kappa/2=1.5\kappa$.
@@
 $n$ $=5,\quad m=384,\quad\delta=10,\quad\epsilon=2^{-40}$
-$\gamma$ $=\left\lceil\frac{128+40+5\times 10+4\times 384}{5}\right\rceil$
-$=\left\lceil\frac{128+40+50+1536}{5}\right\rceil=\lceil 350.8\rceil=351$
+$\gamma$ $=\frac{128+40+5\times 10+4\times 384}{5}$
+$=\frac{128+40+50+1536}{5}=\frac{1754}{5}=350.8\approx 351$
 $H_{\infty}(S)$ $\geq 5\times(351-10)-4\times 384$
@@
 The protocol’s resilience to quantum attacks stems from its layered defense strategy:
+  • 1.
     Structural defense: The XOR operation’s linear algebra properties prevent quantum speedup exploitation
+  • 2.
     Entropic defense: Min-entropy bounds ensure exponential search spaces
+  • 3.
     Cryptographic defense: QROM-based commitments resist quantum collision attacks
-Theorem 9 (Comprehensive Quantum Security).
-For any QPT adversary $\mathscr{A}$ with quantum memory $Q=2^{q}$ qubits, the distinguishing advantage satisfies:
-$\left|\Pr[\mathscr{A}(K)=1]-\Pr[\mathscr{A}(U)=1]\right|$
-$\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\frac{n^{2}}{2^{\lambda}}+2^{q-(n\gamma-(n-1)m)}.$ (34)
+Theorem 5.3 (Comprehensive Quantum Security).
+For any QPT adversary $\mathcal{A}$ with quantum memory $Q=2^{q}$ qubits, the distinguishing advantage satisfies:
+$\left|\Pr[\mathcal{A}(K)=1]-\Pr[\mathcal{A}(U)=1]\right|\leq 2^{-\kappa}+\frac{q^{2}}{2^{m}}+\frac{n^{2}}{2^{\lambda}}+2^{q-(n\gamma-(n-1)m)}.$
 Theorem 9 provides a unified security bound incorporating all quantum attack vectors.
@@
 Our parameterization ensures all terms are $\leq 2^{-128}$ for 128-bit security.
-5-C1 Quantum Collision Resistance Analysis
-The Brassard-Høyer-Tapp (BHT) attack [7] achieves $O(2^{m/3})$ complexity by:
-[…]
-[…]
-$\text{Measure collision with prob. }p\geq c\cdot 2^{-m/3}.$
+5.3.1 Quantum Collision Resistance Analysis
+The Brassard-Hoyer-Tapp (BHT) attack [7] achieves $O(2^{m/3})$ complexity by:
+Phase 1: Create $\frac{\pi}{4}2^{m/3}$ states $|\psi_{i}\rangle=\sum\limits_{x}\alpha_{x}|x\rangle$
+Phase 2: Apply $U_{f}:|x\rangle\rightarrow(-1)^{f(x)}|x\rangle$
+Measure collision with prob. $p\geq c\cdot 2^{-m/3}$.
 Our parameterization $m=3\kappa$ ensures:
-$\text{Expected queries}=\sqrt{\frac{\pi}{4p}}\geq\sqrt{\frac{\pi}{4c}}\cdot 2^{\kappa}\gg 2^{\kappa}.$ (35)
+$\text{Expected queries}=\sqrt{\frac{\pi}{4p}}\geq\sqrt{\frac{\pi}{4c}}\cdot 2^{\kappa}\gg 2^{\kappa}.$
 The commitment structure $c_{i}=\mathcal{H}(s_{i}\parallel\hat{H}_{i})$ forces domain separation, preventing Wagner’s generalized birthday attacks in quantum settings.
 6 Entropy Requirements and Parameter Trade-offs
-6-A Theoretical Parameter Optimization
+6.1 Theoretical Parameter Optimization
 Optimize parameters for security and efficiency ($\kappa=128$, $\delta=10$, $m=384$, $\epsilon=2^{-40}$):
-Table 2 incorporates 16-bit security margins to mitigate estimation uncertainties and unforeseen attacks.
-The margin $H_{\infty}(S)-\kappa$ absorbs potential reductions from quantum side information $S_{0}(\rho_{E})$.
-For $n=4$, $\gamma$ increases to 340 ensuring $H_{\infty}(S)\geq 152>\kappa+24$, satisfying:
-$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)+\zeta$ (36)
-where $\zeta=24$ represents the operational security buffer. This conservative parameterization accounts for possible deviations in entropy estimation and quantum memory effects.
 Table 1: Parameterization for $\kappa=128$-bit Quantum Security
@@
 7 384 359 179 3.15 51
+Table 1 incorporates 16-bit security margins to mitigate estimation uncertainties and unforeseen attacks.
+The margin $H_{\infty}(S)-\kappa$ absorbs potential reductions from quantum side information $S_{0}(\rho_{E})$.
+For $n=4$, $\gamma$ increases to 340 ensuring $H_{\infty}(S)\geq 152>\kappa+24$, satisfying:
+$n(\gamma-\delta)-(n-1)m\geq\kappa+\log(1/\epsilon)+\zeta$
+where $\zeta=24$ represents the operational security buffer. This conservative parameterization accounts for possible deviations in entropy estimation and quantum memory effects.
 Note: Communication cost computed as $\frac{n(n-1)m}{8\times 1024}$ KB
 Design Guidelines:
+  • 1.
     Small $n$: Higher $\gamma$ required, but lower communication
+  • 2.
     Large $n$: Lower $\gamma$ possible, but quadratic communication overhead
+  • 3.
     Balanced: $n=5$ provides optimal tradeoff for 128-bit security
+  • 4.
     Security margin: $H_{\infty}(S)-\kappa$ provides buffer against unforeseen attacks
-6-A1 Quantum-Secure Entropy Estimation
+6.1.1 Quantum-Secure Entropy Estimation
 Conventional min-entropy estimators exhibit bias under quantum sampling. Adopting quantum-proof techniques [33], we bound estimation error:
-Theorem 10 (Quantum Min-Entropy Sampling).
-For $\epsilon>0$ and samples $\mathcal{S}=\{x_{1},\dots,x_{N}\}$ from source $X$:
-$\Pr_{x_{i}\leftarrow X}\left[\hat{H}_{\infty}(X)\geq S_{\infty}(X|E)_{\rho}-\Delta\right]\geq 1-\epsilon$ (37)
+Theorem 6.1 (Quantum Min-Entropy Sampling).
+For $\epsilon>0$ and samples $\mathcal{S}=\{x_{1},\ldots,x_{N}\}$ from source $X$:
+$\Pr_{x_{i}\leftarrow X}\left[\hat{H}_{\infty}(X)\geq S_{\infty}(X|E)_{\rho}-\Delta\right]\geq 1-\epsilon$
 with $\Delta=\log(1/\epsilon)+\frac{1}{2}\log|\mathcal{X}|-H_{2}(X)$. When $N\geq\frac{2}{\epsilon^{2}}\log(2|\mathcal{X}|)$, $\Delta\leq 2\delta$ for $\delta$ in Definition 3.2.
@@
 This adaptivity is crucial for maintaining composable security under quantum side information.
-6-B Quantum Advantage Mitigation
+6.2 Quantum Advantage Mitigation
 To counter quantum speedups:
+  • 1.
     Grover mitigation: Set $\kappa^{\prime}=2\kappa$ for 128-bit quantum security
+  • 2.
     BHT mitigation: Set $m\geq 3\kappa$
+  • 3.
     Quantum memory: Set $\gamma\geq\kappa+\log Q$
+  • 4.
     Error margin: Include $\delta$ buffer for entropy estimation
+  • 5.
     Composability: Use QUC framework for modular security [6]
-6-B1 Quantum Effects on Entropy Estimation
+6.2.1 Quantum Effects on Entropy Estimation
 Conventional min-entropy estimators exhibit vulnerabilities under quantum sampling:
-$\sup_{\rho}\left|\hat{H}_{\infty}^{\text{class}}(X)-S_{\infty}(X|E)_{\rho}\right|\leq\delta+\log(1/\epsilon)$ (38)
+$\sup_{\rho}\left|\hat{H}^{\text{class}}_{\infty}(X)-S_{\infty}(X|E)_{\rho}\right|\leq\delta+\log(1/\epsilon)$
 Our solution employs quantum-proof estimators via:
   1. 1.
-    Quantum-secure randomness extractors: $\text{Ext}:\{0,1\}^{m}\times\{0,1\}^{d}\rightarrow\{0,1\}^{\kappa}$
+    Quantum-secure randomness extractors: Ext : $\{0,1\}^{m}\times\{0,1\}^{d}\rightarrow\{0,1\}^{\kappa}$
   2. 2.
@@
     Min-entropy sampling from quantum sources [15]:
-    $\Pr_{x^{n}\leftarrow X^{n}}\left[\hat{H}_{\infty}(X)\geq H_{\infty}(X)-\Delta\right]\geq 1-\epsilon$ (39)
+    $\Pr_{x^{n}\leftarrow X^{n}}\left[\hat{H}_{\infty}(X)\geq H_{\infty}(X)-\Delta\right]\geq 1-\epsilon$
     where $\Delta=O(\sqrt{n}/|\mathcal{X}|)$.
@@
 7 Comparative Analysis and Protocol Extensions
-7-A Comparative Analysis with Post-Quantum Alternatives
+7.1 Comparative Analysis with Post-Quantum Alternatives
 The quantum security landscape features diverse approaches with fundamentally different security foundations.
 Our protocol’s distinctive information-theoretic security provides unique advantages compared to computational post-quantum solutions.
-As shown in Table 3, while lattice-based schemes like CRYSTALS-Kyber offer efficient $\mathcal{O}(n\kappa)$ communication, their security relies on the unproven hardness of module-LWE problems, which remain vulnerable to unforeseen quantum algorithmic breakthroughs.
-Similarly, isogeny-based schemes like SIKE provide compact key sizes but have suffered devastating cryptanalytic attacks in recent years, demonstrating the fragility of dependency on specific mathematical assumptions.
-TABLE 3: Post-Quantum Key Agreement Protocol Comparison
-Protocol Security Basis Sec. Type Comm. Comp. Hardware Trust Hardness
-Our Protocol Rényi entropy Info-theoretic $\mathcal{O}(n^{2}m)$ $\mathcal{O}(n^{2})$ Classical Honest majority None
-CRYSTALS-Kyber [8] Module-LWE Computational $\mathcal{O}(n\kappa)$ $\mathcal{O}(n\kappa^{2})$ Classical PKI MLWE
-SPHINCS+ [28] Hash functions QROM $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa)$ Classical None QROM
-SIKE [11] Isogenies Computational $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa^{4})$ Classical PKI SSI
-QKD [25] Quantum physics Info-theoretic $\mathcal{O}(\kappa)$ $\mathcal{O}(\kappa)$ Quantum Auth. chan. None
-Sec. Type: Security Type; Comm.: Communication Complexity; Comp.: Computational Complexity;
-Hardware: Hardware Requirements; Trust: Trust Assumptions; Hardness: Hardness Assumptions.
-Auth. chan.: Authenticated channels; PKI: Public Key Infrastructure.
+As shown in Table 2, while lattice-based schemes like CRYSTALS-Kyber [8] offer efficient $\mathcal{O}(n\kappa)$ communication, their security relies on the unproven hardness of module-LWE problems, which remain vulnerable to unforeseen quantum algorithmic breakthroughs.
+Similarly, isogeny-based schemes like SIKE [11] provide compact key sizes but have suffered devastating cryptanalytic attacks in recent years, demonstrating the fragility of dependency on specific mathematical assumptions.
 Quantum key distribution (QKD) shares our information-theoretic security properties but requires specialized quantum communication hardware and authenticated classical channels.
@@
 For $\kappa=128$-bit security with $n=5$ parties, we achieve $H_{\infty}(K)\geq 169$ bits, providing a security buffer against unforeseen attacks.
-7-B Protocol Extensions to Secure Applications
+Table 2: Comparison with Post-Quantum Alternatives
+Approach Security Basis Advantages Limitations
+Our Protocol Information-theoretic Unconditional security, Quantum resistance, Fault tolerance Quadratic communication, Entropy source requirement
+Lattice-based (e.g., Kyber) Computational (LWE) Efficient, Standardized Vulnerable to quantum algorithmic advances
+Code-based Computational (Decoding) Mature theory, Conservative security Large key sizes, Not efficient
+QKD Information-theoretic Proven security, Commercial availability Requires quantum channels, Distance limitations
+7.2 Protocol Extensions to Secure Applications
 The core protocol naturally extends to several high-impact applications through novel cryptographic frameworks:
-1.
-Secure Multiparty Computation for Linear Functions: The protocol directly supports privacy-preserving computation of linear functions $f(s_{1},\dots,s_{n})=\sum c_{i}s_{i}$ without revealing individual inputs.
+  1. 1.
+    Secure Multiparty Computation for Linear Functions: The protocol directly supports privacy-preserving computation of linear functions $f(s_{1},\ldots,s_{n})=\sum c_{i}s_{i}$ without revealing individual inputs.
 This enables:
+    • (a)
       Federated learning: Secure aggregation of model updates while preserving data privacy
+    • (b)
       Financial auditing: Cross-institutional fraud detection with confidential inputs
+    • (c)
       Supply chain optimization: Collaborative logistics planning with proprietary data protection
-The derived key $K_{f}=\mathcal{H}(f(s_{1},\dots,s_{n}))$ maintains the min-entropy guarantee $H_{\infty}(K_{f})\geq\kappa$ when coefficients $c_{i}$ are properly constrained.
-2. Quantum-Hybrid Security Architecture: Integrating with quantum key distribution creates a defense-in-depth architecture:
-$\text{Hybrid Key}=\mathcal{H}(K_{\text{QKD}}\oplus K_{\text{Entropy}})$ (40)
+    The derived key $K_{f}=\mathcal{H}(f(s_{1},\ldots,s_{n}))$ maintains the min-entropy guarantee $H_{\infty}(K_{f})\geq\kappa$ when coefficients $c_{i}$ are properly constrained.
+  2. 2.
+    Quantum-Hybrid Security Architecture: Integrating with quantum key distribution creates a defense-in-depth architecture:
+    $\text{Hybrid Key}=\mathcal{H}(K_{\text{QKD}}\oplus K_{\text{Entropy}})$
     This combines the active security of QKD with the fault tolerance of our entropy protocol, creating a quantum-resistant solution suitable for critical infrastructure. The hybrid approach provides:
+    • (a)
       Enhanced active security through QKD’s authentication mechanisms
+    • (b)
       Fault tolerance against malicious participants via secret sharing
+    • (c)
       Defense in depth where compromise of one system doesn’t break overall security
-3. Post-Quantum Authentication Framework: We develop an entropy-based message authentication code (MAC):
-$\text{MAC}(k,m,s)=\mathcal{H}(k\oplus\mathcal{H}(m\|s))$ (41)
+  3. 3.
+    Post-Quantum Authentication Framework: We develop an entropy-based message authentication code (MAC):
+    $\text{MAC}(k,m,s)=\mathcal{H}(k\oplus\mathcal{H}(m\|s))$
     leveraging the same entropy sources used in key agreement.
-This provides quantum-resistant authentication with security bound $\Pr[\text{forge}]\leq\frac{q^{2}}{2^{\gamma}}+\frac{(q+1)^{2}}{2^{\lambda}}$, integrating seamlessly with our key establishment protocol.
-7-C Future Research Directions
+This provides quantum-resistant authentication with security bound $\Pr[\text{forge}]\leq\frac{q^{2}}{2\gamma}+\frac{(q+1)^{2}}{2\lambda}$, integrating seamlessly with our key establishment protocol.
+7.3 Future Research Directions
 Four high-impact research directions emerge from this work:
-1) Lightweight Implementations for IoT: Developing optimized implementations for resource-constrained devices presents significant challenges.
+  1. 1.
+    Lightweight Implementations for IoT: Developing optimized implementations for resource-constrained devices presents significant challenges.
 The quadratic communication complexity $\mathcal{O}(n^{2}m)$ becomes problematic for large $n$, requiring compression techniques for shares and commitments.
 Research should explore:
+    • (a)
       Hierarchical secret sharing to reduce communication
+    • (b)
       Efficient entropy estimation with limited samples
+    • (c)
       Hardware acceleration for polynomial operations
-2) Blockchain Integration for Randomness Beacons: The protocol can power decentralized randomness generation for blockchain consensus:
-$K_{\text{epoch}}=\mathcal{H}\left(\bigoplus_{i=1}^{n}\mathcal{H}(s_{i}\|\text{block}_{h-1})\right)$ (42)
+  2. 2.
+    Blockchain Integration for Randomness Beacons: The protocol can power decentralized randomness generation for blockchain consensus:
+    $K_{\text{epoch}}=\mathcal{H}\left(\bigoplus_{i=1}^{n}\mathcal{H}(s_{i}\|\text{block}_{h-1})\right)$
     providing verifiable randomness with min-entropy bound $H_{\infty}(K_{\text{epoch}}|\mathcal{B})\geq\gamma-\log(n|\mathcal{B}|)$. This enables:
+    • (a)
       Bias-resistant consensus protocols
+    • (b)
       Fair NFT minting and airdrops
+    • (c)
       Transparent on-chain gambling
-3) Post-Quantum Cryptography Standardization: Our work provides foundations for standardization efforts in information-theoretic PQC. Key initiatives include:
+  3. 3.
+    Post-Quantum Cryptography Standardization: Our work provides foundations for standardization efforts in information-theoretic PQC. Key initiatives include:
+    • (a)
       Formal security proofs using quantum proof assistants
+    • (b)
       Parameterization guidelines for different security levels
-  • Implementation testing frameworks Side-channel resistance certification
-4) Entropy Source Diversity and Robustness: Future work must address practical challenges in entropy source management:
+    • (c)
+      Implementation testing frameworks
+    • (d)
+      Side-channel resistance certification
+  4. 4.
+    Entropy Source Diversity and Robustness: Future work must address practical challenges in entropy source management:
+    • (a)
       Security against adversarial entropy sources
+    • (b)
       Cross-source min-entropy estimation techniques
+    • (c)
       Quantum-resistant entropy pooling methods
+    • (d)
       Continuous entropy validation during operation
@@
 8 Conclusion
-We have presented a theoretical framework for post-quantum key agreement based on Rényi entropy.
+We have presented a theoretical framework for post-quantum key agreement based on Renyi entropy.
 The enhanced protocol addresses vulnerabilities in prior constructions through innovations: (1) a confidentiality-preserving verification mechanism using distributed polynomial commitments, (2) provably non-negative min-entropy bounds for XOR composition, and (3) composable security proofs in the quantum universal composability model.
 The protocol’s security rests on three mathematical pillars derived from quantum information theory:
-Entropy Preservation: $H_{\infty}(S)$ $\geq n\gamma-(n-1)m$ (43)
-Verification Security: $\Pr[\text{bypass}]$ $\leq 2^{-\kappa}$ (44)
-Composable Security: $\delta_{\text{QUC}}$ $\leq\operatorname{negl}(\kappa)$ (45)
-Equations (43)-(45) form an integrated security framework that resists quantum attacks through fundamental information-theoretic principles rather than computational assumptions.
+Entropy Preservation: $\quad H_{\infty}(S)\geq n\gamma-(n-1)m$
+Verification Security: $\quad\Pr[\text{bypass}]\leq 2^{-\kappa}$
+Composable Security: $\quad\delta_{\text{QUC}}\leq\text{negl}(\kappa)$
+These equations form an integrated security framework that resists quantum attacks through fundamental information-theoretic principles rather than computational assumptions.
 For 128-bit security, our parameter optimization yields:
@@
 Key advantages include:
+  • 1.
     Information-theoretic security without reliance on hardness assumptions
+  • 2.
     Resistance to quantum algorithm breakthroughs
+  • 3.
     Fault tolerance via secret sharing
+  • 4.
     Extensibility to multiparty computation and hybrid systems
@@
 By leveraging information-theoretic principles, this work establishes a paradigm for long-term cryptographic security in the quantum era.
-Biography Section
-Ruopengyu Xu received his B.S degree in network engineering from Software Engineering Institute Gungzhou in June 2025.
-He is currently an independent scientific researcher.
-Recently, he has published a paper in https://arxiv.org/abs/2506.07640.
-At the same time, he is also preparing his postgraduate plan.
-His research interests are in Quantum cryptography, Information Security, Number Theory and Public key cryptosystem so on.
-Chenglian Liu received his B.S.
-degree in information management from National Union University in 1992, and his M.Sc.
-degree in National Defense from National Defense University in 2004.
-He was a research fellow in the Department of Automation Engineering at Technological Educational Institution of Sterea Hellas from 2017 to 2019, and a visiting scholar at Universidad del Pacifico from 2020 to 2022.
-He is currently an associate professor at the School of Electrical and Computer Engineering, Nanfang University, Guangzhou, China.
-His main research focuses on Information Security, Cryptanalysis, Fuzzy Sets, and Goldbach’s Conjecture, among other topics.
+Appendix A Proofs of Technical Lemmas
+This appendix contains detailed proofs of selected technical results from the main text.
+A.1 Proof of Theorem 4 (Complete Version)
+The complete proof of the min-entropy preservation theorem involves careful analysis of the convolution properties of probability distributions over finite fields…